### v1

# Unsupervised Episode Detection for Large-Scale News Events

Abstract.
Episodic structures are inherently interpretable and adaptable to evolving large-scale key events. However, state-of-the-art automatic event detection methods overlook event episodes and, therefore, struggle with these crucial characteristics. This paper introduces a novel task, episode detection, aimed at identifying episodes from a news corpus containing key event articles. An episode describes a cohesive cluster of core entities (e.g., “protesters”, “police”) performing actions at a specific time and location. Furthermore, an episode is a significant part of a larger group of episodes under a particular key event. Automatically detecting episodes is challenging because, unlike key events and atomic actions, we cannot rely on explicit mentions of times and locations to distinguish between episodes or use semantic similarity to merge inconsistent episode co-references. To address these challenges, we introduce EpiMine, an unsupervised episode detection framework that (1) automatically identifies the most salient, key-event-relevant terms and segments, (2) determines candidate episodes in an article based on natural episodic partitions estimated through shifts in discriminative term combinations, and (3) refines and forms final episode clusters using large language model-based reasoning on the candidate episodes. We construct three diverse, real-world event datasets annotated at the episode level. EpiMine outperforms all baselines on these datasets by an average 59.2% increase across all metrics.
Unsupervised Episode Detection, Event Detection, Segment Classification, Large Language Models, Key Term Extraction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Information extractionccs: Information systems Data mining
Figure 1. Example of an event structure hierarchy. We take a key event corpus as input and aim to detect each episode node with a corresponding cluster of relevant text segments.
1. Introduction
Given the saturation of real-time news accessible at our fingertips, reading and processing the critical information of a key event has become an increasingly daunting challenge. Consequently, research on automatic textual event detection has recently attempted to integrate the manner in which humans neurologically perceive/store events into textual event detection methods. Specifically, neuroscientists studying event representations in human memory find that events are stored in a top-to-bottom hierarchy, as demonstrated in Figure 1. The deeper the hierarchical event level, the more fine-grained its corresponding text granularity (Zhang et al., 2022): we consider a theme as corpus-level (all articles discussing the 2019 Hong Kong Protests), key event as document-level (an article typically discusses a full one to two day key event), episode as segment-level, and atomic action as sentence or phrase-level.
Furthermore, neurological research (Baldassano et al., 2017; Khemlani et al., 2015) indicates that events are encoded into memory as episodic structures. Representing events as discrete episodes helps us piece together a coherent narrative by considering the sequence of actions, reactions, and developments over time. This empowers several downstream tasks (e.g., event schema generation, event prediction), which may benefit from insights into the causes and consequences of events that might be missed with a more generalized analysis. Despite its strengths, existing automatic event extraction works fail to consider the episode-level.
For example, key event detection specifically seeks to output “a set of thematically coherent documents” for each key event (Zhang et al., 2022; Liu et al., 2023). However, it is challenging to manually parse through a large cluster of relevant articles in order to gain an efficient and compact understanding of a given event. To address the lack of interpretability of such article clusters, the adjacent task of timeline summarization (Steen and Markert, 2019; Li et al., 2021b; Gholipour Ghalandari and Ifrim, 2020; Chen et al., 2023) aims to identify the dates and a compact summary for each key event. However, high-level timelines are typically applicable for historical themes and thus unrealistic for currently evolving key events where fine-grained timeline summarization is more suitable. Hence, event chain mining (Jiao et al., 2023) attempts to address this by mining a series of temporally-ordered atomic actions at the phrase-level; however, the granularity level is often too fine-grained to properly represent a large-scale key event. Thus, we aim to tackle the novel task of episode detection to pave the way for a more effective event representation.
Episode detection aims to detect episodes from a news corpus containing key event articles. An episode can be described as a cohesive cluster of subjects performing actions at a certain time and location, occurring as part of a larger sequence of episodes under a specific key event. Episode detection introduces a unique set of challenges, which we address using our novel framework, EpiMine. EpiMine is an unsupervised episode detection framework that automatically detects meaningful episodic events and their corresponding text segments in a large key event corpus, all without any level of human supervision or labeled training data. EpiMine is comprised of the following components: (1) discriminative co-occurrence detection, (2) episode partitioning, (3) candidate episode estimation, and (4) episode-segment classification. Collectively, they tackle the unique challenges of episode detection, detailed below:
Challenge 1: Journalists do not timestamp episodes. Key event detection partitions a thematic corpus into document-level clusters by heavily relying on explicit temporal features, like publication dates (Zhang et al., 2022). For example, an article primarily discusses one key event, which can then be roughly mapped to the article’s publication date. However, this assumption fails at the episode-level, where there is no guarantee to have a distinct timestamp associated with each text segment that discusses a new episode. Fortunately, we can take advantage of the idea that journalists naturally partition news articles by sequentially discussing distinct episodes: {adjustwidth}1em0em Example: An article likely completes its discussion of the episode A, protesters storming the Legislaive Council, before episode B, “protesters vandalized the Legislative Chamber” (Figure 3). Hence, in order to partition articles into distinct episode segments, EpiMine must identify whether or not two consecutive segments are discussing the same or different episodes, which brings us to our next challenge.
Challenge 2: Episodes contain semantically diverse actions. Each episode features a set of unique atomic actions, which we can utilize for determining whether or not two segments discuss the same episode. However, for clustering actions, existing methods (Jiao et al., 2023) rely heavily on semantic similarity. This not realistic for episode-segment clustering: {adjustwidth}1em0em Example:protesters spray-painted slogans” and “they unfurled the colonial-era flag” will fall under the same episode, but are semantically different and unlikely to be clustered. Alternatively, we can identify salient terms that reflect the same episode (“barriers” and “shoved” are unique to Episode A; “defaced” and “walls” for Episode B), by exploiting corpus-level signals. Specifically, if we frequently see “defaced” mentioned together with “walls” (or their respective synonyms) and not with other terms (and vice-versa), then we consider them a discriminative co-occurrence. Consequently, when such co-occurrences remain consistent across text segments, this indicates the same episode being discussed. Conversely, if a sufficient shift in the combination of terms occurs, then this indicates that a different episode is being discussed.
Challenge 3: Articles often do not feature all episodes. Given the real-time nature of reporting, an article may feature only a subset or none of ground-truth episodes from a multi-day key event (e.g., focusing on a high-level analysis of the key event or mentioning other related key events). To minimize this noise, EpiMine seeks to identify the set of articles which maximizes the quantity and quality of potential episodes. It then merges any article partitions across these articles which likely discuss the same episode and employs a large language model (LLM) to provide a more fluent interpretation of the candidate episodes, accounting for the episode’s core entity, actions, object, location, and time period. This allows EpiMine to finally map the remaining non-salient article segments to these episodes, pruning any candidates which are not sufficiently supported by the remaining articles. We summarize our core contributions:
  1. (1)
    We introduce the novel task of episode detection, which takes in a key-event corpus and outputs multiple detected episodes and their related segments extracted from the corpus.
  2. (2)
    We propose a novel unsupervised episode detection method, EpiMine, which exploits discriminative term co-occurrences to estimate candidate episode partitions from top articles. It refines these candidate episodes using LLM-based reasoning to output a comprehensive and diverse set of episodes.
  3. (3)
    We construct three novel datasets, reflecting a diverse set of real-world themes and thirty global key events, as no large-scale key event-specific news corpus exists for this task where the key events are guaranteed to contain distinguishable episodes.
  4. (4)
    We compare EpiMine with five other baselines through an extensive set of experiments and case studies performed on our real-world datasets, demonstrating that it outperforms all baselines by, on average, a 59.2% increase across all metrics.
Reproducibility: We provide our dataset and source code [^1] https://github.com/pkargupta/epimine to facilitate further studies.
2. Related Works
2.1. Event Extraction
Understanding events in unstructured texts is crucial, leading to substantial research in event detection and extraction (Liu et al., 2018a; Du and Cardie, 2020; Li et al., 2021a; Lu et al., 2021; Qi et al., 2022; Jiao et al., 2022). Existing work, such as Han et al. (2019); Wang et al. (2020); Ahmad et al. (2021), have focused on event relation extraction and prediction, while salient event identification remains a significant area of study (Liu et al., 2018b; Jindal et al., 2020; Wilmot and Keller, 2021). Furthermore, event process understanding has garnered recent attention (Zhang et al., 2020; Chen et al., 2020). However, these studies typically require costly expert annotations. Some efforts have sought to mitigate this challenge through unsupervised approaches (Weber et al., 2018; Li et al., 2020). The pioneering work (Chambers and Jurafsky, 2008) introduced event chains as a novel representation of structured news knowledge. In the realm of biological reading comprehension, Berant et al. (2014) extracted events and their relationships in biological processes. More recently, Zhang et al. (2021) utilized event graphs constructed from large corpora to derive event chains for narrative understanding. Despite these advancements, most existing studies concentrate on phrase-level or sentence-level events (analogous to actions in this paper), neglecting the hierarchical structures of events at varying granularities within textual data. Additionally, these approaches often rely on human-curated event ontologies, are restricted to predefined event types, and fail to capture events from open-domain texts. Although key event extraction has been explored in some studies (Zhang et al., 2022; Jiao et al., 2023), the automatic detection of episodes remains under-explored and critically important.
2.2. Timeline Summarization
Timeline summarization (TLS) automatically identifies key dates of major events and provide concise descriptions of occurrences on those dates. Previous approaches to TLS have predominantly focused on extractive methods. For instance, Nguyen et al. (2014) built thematic timelines for general-domain topics by selecting and ranking events relevant to an input query. Martschat and Markert (2018) introduced a temporally sensitive submodularity framework, designing objective functions and constraints that model the temporal dimension inherent in timeline summarization. Steen and Markert (2019) were the first to propose an abstractive timeline summarization system. Their approach clusters sentences likely describing the same event and uses multi-sentence compression to generate summary sentences for each cluster. More recently, Li et al. (2021b) represented news articles as event-graphs, transforming the summarization task into compressing the graph to its salient sub-graph. Additionally, You et al. (2022) presented a joint learning-based heterogeneous graph attention network, integrating event detection into a unified framework to eliminate redundant sentences. While these methods involve key event grouping and understanding, they typically apply to historical themes, producing higher-level timelines. However, these approaches may not be practical for ongoing news events, where a detailed timeline summary (distinguishing between episodes) is more appropriate.
2.3. Topic Discovery
The goal of topic discovery is to group news articles by their themes, an area of study dating back to the early days of automated event theme identification from textual data (Allan et al., 1998). One approach (Zhou et al., 2015; Beykikhoshk et al., 2018) treats this task as a topic modeling issue, using LDA (Blei et al., 2003) or its enhanced versions. Another line of research (Sayyadi et al., 2009; Sayyadi and Raschid, 2013) constructs a graph based on keyword frequency and identifies themes from keyword clusters, mapping these themes to documents via keyword-driven feature similarities. However, these techniques do not consider temporal information, limiting them to discerning different thematic events. To address this, some studies (Fung et al., 2005; He et al., 2007) introduce tracking keyword spikes to pinpoint events. Specifically, (Zhao et al., 2012) uses burst detection (Kleinberg, 2004) to identify and utilize bursty terms for document representation and clustering. Building on this, (Ge et al., 2016) creates a burst information network with recognized bursts and their attributes for clustering. Finally, (Yoon et al., 2023) proposes unsupervised stream-based story discovery, which computes article embeddings their shared temporal themes. Overall, these methods solely focus on the key-event or atomic action level.
Figure 2. We detail the overall framework of EpiMine.
3. Preliminaries
3.1. Problem Definition
Definition 3.1 (Episode).
An episode $E_{i}$ is one of a partially ordered sequence of subevents, $\{E_{1},\ldots,E_{i},\ldots E_{k}\}$, of a key (major) event $E$, where typically $2\geq k\geq 20$, and $E_{i}$ does not overlap with $E_{j}$ if $i\neq j$. Actions in the episode $E_{i}$ can be either semantically similar or diverse, but typically have relatively tight time, location, and/or thematic ( entities, actions, objects) proximity between them.
EpiMine aims to extract episodes from a news corpus, where an episode occurs as a significant component of a larger group of episodes that fall under a specific key event. For instance, in Figure 1, without knowing Episode #1, “Protesters stormed the Legislative Council Complex”, readers would not have fully understood Episode #3, “Police dispersed protesters”. Hence, episodes help us understand the overall key event and are especially useful for events that are currently evolving, where finer context is required for sufficiently understanding them.
Definition 3.2 (Episode Detection).
Given a corpus $\mathcal{D}$ about one key event, where each document $d\in\mathcal{D}$ is a news article, the goal of the task is to obtain the set of text segment clusters =$\mathcal{E}=\{E_{1},E_{2},\ldots,E_{K}\}$. Each episode cluster $E_{i}\subset\mathcal{S}=\{s_{1}^{1},s_{2}^{1},\ldots,s_{|d|}^{|\mathcal{D}|}\}$, where $\mathcal{S}$ contains all of the text segments identified in each document $d\in\mathcal{D}$, and every two clusters do not have overlapping text segments (i.e., $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
It is important to note that the number of episodes $K$ is not known in advance and oftentimes, a news article segment may discuss either episodes of a different key event (e.g., an episode with similar aspects that occurred in a different historical key event) or multiple episodes of the current key event. Nonetheless, our goal is to detect the most relevant episodes to the current key event at hand and consequently mine the most distinctive text segments for each of these (hence our constraint of $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
3.2. Key Event News Article Pre-Processing
Given that the expected output for the episode detection task is a cluster of text segments, we first must segment each key event news article. We would like to ideally preserve both the primary aspects (e.g., core entities and their actions) and peripheral aspects (e.g., reactions to a core entity’s action) relevant to that episode, which may be helpful for cross-document episode co-reference resolution. In order to do this, we utilize the text segmentation method, C99 (Choi, 2000). Furthermore, in order to assist with the cohesiveness of the segment, we employ entity co-reference resolution before performing segmentation, which assists with retaining the context across text segments (“They surrounded the legislative building […]” $\rightarrow$ “The protesters surrounded the legislative building […]”). In the following section, we present our core methodology given these text segments (in their raw form, without co-references resolved) and their source articles as the primary inputs.
4. Methodology
In order to tackle the episode detection task, we propose a novel unsupervised framework, EpiMine. As shown in Figure 2, EpiMine consists of the following four core components: (1) episode indicative term mining, which identifies combinations of salient terms that are likely to discriminatively co-occur within an episode and not across episodes; (2) episode partitioning, which partitions each article into approximate isolated episodes based on consecutive shifts in the term co-occurrence distribution, (3) candidate episode estimation, which clusters the top partitions into candidate episodes and utilizes LLM-based reasoning to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters.
4.1. Episode Indicative Term Mining
Figure 3. Natural partition between two episodes within a key event article. An episode’s discriminative terms are bolded, while salient but non-discriminative terms are underlined.
In the absence of supervision, our goal is to identify potential candidates for episodes. Episodes are often described in relation to each other and usually lack specific timestamps or locations consistently mentioned within their segments. For example, the phrase “police dispersed protesters” may not have a precise timestamp because it is a response to “protesters stormed the Legislative Council Complex,” and some journalists may consider the implicit ordering adequate.
Additionally, the same episode can be described using different entities and actions, as journalists may highlight different perspectives when reporting an episode. For example, both “protesters shoved against the barricades” and “the police used pepper-spray on the protesters” describe the episode “protesters stormed the Legislative Council Complex”. However, they are semantically different, focusing on different core entities and actions. Therefore, we cannot depend on a consistent subject-action-object triple or an explicit time and location associated with each episode in the articles.
Salient Terms.
To circumvent this challenge, we exploit the idea that journalists naturally partition news articles according to episodes, forming episode fragments. For example, as shown in Figure 3, an article will likely complete its discussion of episode #1, “Protesters stormed the Legislative Council Complex” (red), before fully shifting to discussing episode #2, “protesters vandalized the Legislative Chamber” (blue). Across these episode fragments, certain salient terms are featured (e.g., protesters, legislative, vandalizing, podium). We modify the idea of event salience from (Jiao et al., 2023) for the task of episode detection:
Definition 4.1 (Salience).
A term is salient if it is (1) distinct and significant to understanding a given key event, as well as (2) frequently found in a key event’s segments and infrequently in other background/general articles.
We define the salience score of a term $w_{i}$ within segment $s$ as the following function, where $freq(w_{i})$ is the number of key event segments that $w_{i}$ is contained in, $N_{bg}$ is the number of news articles in the background corpus we construct (using general New York Times articles), and $bgf(w_{i})$ is the number of background articles that $w_{i}$ is present in.
(1) $\textbf{Salience}(w_{i})=\left(1+\log^{2}\left(freq(w_{i})\right)\right)\log\left(\frac{N_{bg}}{bgf(w_{i})}\right)$
Stop words and infrequent terms ($freq(w_{i})<5$) are assigned a salience score of $-1$. A key event’s set of salient terms $T$ is comprised of the terms with a salience score above the mean salience across the entire vocabulary. In the case of infrequent synonyms used by a journalist as a stylistic choice (e.g., “demonstrations”, “marches”), we expand $T$ with terms that are similar (cosine-similarity) to their static word representations (average of its contextualized word embeddings across entire key event corpus).
Discriminative Co-occurrence.
Now, each text segment within the key event corpus has a corresponding set of salient terms identified. In Figure 3, we can see that the first episode fragment, and Episode #1 in general, features a combination of similar terms, such as “protesters”, “barrier”, and “breach”. Likewise, the second episode may include a combination of terms similar to “protesters”, “spray-painting”, and/or “flag”. Hence, we introduce the significance of co-occurrence for episode detection. Despite some journalists choosing to only describe the protesters spray-painting while others focus on the protesters draping the colonial-era flag, we must be able to recognize that their respective salient terms are likely to co-occur within the same episode.
However, we make a novel distinction between a co-occurrence and a discriminative co-occurrence. A salient term $a$ (e.g., “protesters”) may often co-occur with another salient term $b$ (“spray-painting”) within an episode. On the other hand, if $a$ frequently co-occurs with many other terms in various episodes (“protesters broke”), $a$ and its co-occurrences are less useful for distinguishing episodes. Thus, $(a,b)$ is not a discriminative co-occurrence.
Definition 4.2 (Discriminative Co-occurrence).
A pair of terms $(a,b)$ discriminatively co-occur if (1) they frequently appear together in episode $E_{i}$, and (2) neither $a$ nor $b$ appear as frequently with other terms $w$ in other episodes $E_{\notin i}$.
We compute the discriminative occurrence $d$ between salient term pair $(a,b)$ using the following equation:
(2) $\begin{split}\textbf{d}(a,b)=&\log\left(\frac{freq(a,b)}{\max(\bar{f_{a}},\bar{f_{b}})}\right)\times\log\left(\frac{|T|}{max(|F_{a}|,|F_{b}|)}\right),\\ \text{where }&\bar{f_{a}}=\frac{1}{|T|}\sum\limits_{\forall w_{i}\in T}freq(a,w_{i})\text{, and}\\ &F_{a}=\{freq(a,w_{i})>1\text{ }\forall\text{ }w_{i}\in T\}\end{split}$
The first $\log$ term ensures that the pair’s co-occurrence ($freq(a,b)$) is statistically significant ($\geq$ the max of $a$ and $b$’s mean vocabulary-wide co-occurrence respectively). The second $\log$ term ensures that the salient term pair is a discriminative matching, penalizing pairings where either $a$ or $b$ co-occurs with a large proportion of the salient term set $T$). For example, co-occurrences with “protesters” are not discriminative because “protesters” is a core entity in all episodes and thus frequently co-occurs with many terms in $T$. In contrast, (“slogans”, “flags”) is a discriminative co-occurrence since both terms frequently appear together in segments discussing episode #2 and rarely co-occur with other terms $w_{i}\in T$. If $a$ and $b$ are the same term or close synonyms (determined by statistically significant semantic similarity), they have maximum co-occurrence. By leveraging multiple articles in a large key event corpus, we have sufficient statistical support to ensure our output reflects the average realistic reporting of the key event and its episodes.
4.2. Episode Partitioning
With the ability to identify discriminative co-occurrences, we can use a key transitive property to resolve episode co-references within and across articles, where not all combinations of an episode’s discriminative terms explicitly co-occur:
If $(a,b)$ and $(b,c)$ are both discriminative co-occurrences, then $(a,c)$ is also likely to be a discriminative co-occurrence.
To illustrate this, we have the following text segment excerpts of a news article (the salient and discriminative terms are italicized):
  1. (1)
    Protesters defaced the emblem of Hong Kong, spray-painted slogans, and unfurled the colonial-era flag.
  2. (2)
    The portrait of LegCo president Andrew Leung was defaced.
  3. (3)
    A slogan on the wall reads: “The government forced us to revolt”.
  4. (4)
    Police said at least 13 people had been arrested on suspicion of involvement in the pro-democracy protest.
We can naturally see that segments 1-3 all discuss the “protesters vandalized the Legislative Chamber” episode, while segment 4 discusses the “police dispersed protesters” episode. We can systematically replicate this partitioning process by considering the discriminative co-occurrence score between all pairwise combinations of terms from segments ($i-1$) and ($i$). If the average discriminative co-occurrence and static semantic similarity between each term $a$ from ($i-1$) and $b$ from ($i$) is statistically significant ($\geq\mu_{d}-\sigma_{d}$) for that specific article $d$ (e.g., notably (slogans, defaced) for segments 1-3), we hypothesize that the same episode is being discussed and merge them into one episode fragment. If not (e.g., (slogans, arrested) for segments 3-4), this indicates that a different episode is being discussed, and we partition them into two episode fragments. Further implementation details are provided in Appendix B.
4.3. Candidate Episode Estimation
After partitioning each article into episode fragments, we can identify the set of articles that maximizes the quantity and quality of potential episodes. This step is crucial as it enables EpiMine to minimize noise from articles that may discuss only a subset or none of the ground-truth episodes (e.g., focusing on high-level analyses or other relevant key events). Simultaneously, it improves our method’s efficiency by not relying on the entire key event corpus. We compute the per-article rank by multiplying two metrics:
  1. (1)
    Quality of episode fragments: A top article should primarily consist of episode fragments containing salient terms that discriminatively co-occur. This reduces the rank of general fragments which summarize/analyze the event. We use the average of each episode fragment’s mean inner-discriminative co-occurrence (across all pairwise combinations of a fragment’s salient terms).
  2. (2)
    Quantity of episode fragments: A top article should ideally contain the highest number of ground-truth episodes. Therefore, we take the $\log$ of the number of episode fragments in the article.
After ranking all articles, we select the top $\delta$to serve as the basis for candidate episode estimation. To estimate the candidate episodes, we must first resolve potential co-references to the same episode across these top articles. We apply agglomerative clustering (Murtagh and Contreras, 2012) to the top episode fragments using a pre-computed distance matrix. The distance between two fragments (inversed) is calculated using the same discriminative and static semantic similarity score used in Section 4.2). We then prune clusters that do not contain a statistically significant number of episode fragments.
Finally, we use an LLM to resolve (1) the lack of time and location stamps in an episode fragment and (2) semantic inconsistency between the fragments within a cluster. The LLM leverages its reasoning capabilities to estimate the candidate episodes by identifying the core attributes of each unique cluster (entities, actions, objects, location, and time). It outputs the episode attributes, relevant keywords (assists with extraction), and top extracted text segments based on the clusters (prompt & example in Appendix C).
4.4. Episode-Segment Classification
With these core summaries of the episode clusters, we obtain a generalized description of each candidate episode. This allows us to compute a simple episode representation for assigning an episode and confidence score to each input segment.
Episode & Segment Representations.
We compute a candidate episode representation by encoding both its LLM core attributes and extracted segments using SentenceTransformers (ST) (Reimers and Gurevych, 2019). Following extremely weakly supervised text classification works, such as (Wang et al., 2021; Kargupta et al., 2023), we take the harmonic mean of these representations, as the latter extracted segments are likely not as significant as the earlier extractions and core attributes. We similarly encode all input segments with the same ST model.
Episode-Segment Confidence Estimation
Following recent work on extremely weakly supervised text classification (Kargupta et al., 2023), we compute a confidence score specific to the episode-segment classification task. Simply mapping a text segment to its top episode (based on highest cosine similarity) overlooks the presence of episode-irrelevant text segments and fails to address segments discussing multiple episodes (e.g., a journalist’s key event summary at the beginning of an article). We aim to avoid assigning such segments to any cluster, ensuring no overlap between episode clusters, as discussed in Section 3.1.
Thus, we compute segment $s_{i}$’s cosine similarity to its top two episodes ($e_{i}^{0}$ and $e_{i}^{1}$). A larger gap ($e_{i}^{0}-e_{i}^{1}$) between the two indicates higher confidence in mapping $s_{i}$ to its top episode. We normalize each gap by the sum of all segment episode gaps within the entire corpus, ensuring the confidence is relative to the key event:
(3) $s_{i,\text{confidence}}=\frac{e_{i}^{0}-e_{i}^{1}}{\sum_{l=1}^{|\mathcal{S}|}(e^{0}_{l}-e^{1}_{l})}$
After computing each segment’s confidence in its top episode, we assign only those segments with a statistically significant confidence level to their respective episode clusters $E$. We then prune candidate episode clusters with $\leq 0$ assigned segments, resulting in our final detected episodes and their clusters, $\mathcal{E}$.
5. Experimental Setup
For implementing EpiMine, we use the following hyperparameters across all datasets: $\delta=25\%$, $sim\_thresh=0.75$. All other hyperparameters are set to their respective default values. To determine statistical significance, we check for whether or not a value is $\geq\mu-\sigma$. For our word representations, we use bert-base-uncased. For our sentence representations, we use all-mpnet-base-v2. We choose to use Claude-2.1222claude.ai/ for fluent candidate episode estimation due to its strong structured JSON/XML input and output formatting abilities. However, this proprietary model can be replaced with any open-source model as EpiMine is model-agnostic. We use only one NVIDIA GeForce GTX 1080 for all experiments.
Table 1. Statistics of our collected datasets. The numbers are averaged per key event.
Theme # docs # episodes # segments
Terrorism/Attacks 32.2 5.9 290.3
Natural Disasters 36.2 7.4 324.6
Political Events 70.2 7.5 667.7
Table 2. Results averaged across each theme, including the mean # of episodes EpiMine identifies per theme. Results are computed on each key event corpus using the top-5 documents for each detected episode. Due to variance in LLM generation, we run it 10 times and report the average of each measure. Bold values denote the top method; $\dagger$ denotes the second-best method.
Terrorism (avg. # of eps = 5.36) Natural Disasters (avg. # of eps = 7.4) Politics (avg. # of eps = 7.5)
Methods 5-prec 5-recall 5-F1 5-prec 5-recall 5-F1 5-prec 5-recall 5-F1
EMiner 0.0864 0.0025 0.0048 0.1037 0.0019 0.0037 0.0866 0.0016 0.0032
K-means 0.2123 0.2123 0.2123 0.2785 0.2847 0.2814 0.1604 0.1604 0.1604
K-means + LLM 0.2858 0.1404 0.1826 0.3740 0.1658 0.22 0.2718 0.1736 0.1825
EvMine 0.2303 0.1502 0.1745 0.2815 0.0802 0.1225 0.0536 0.04 0.0458
EvMine + LLM 0.3788 0.1570 0.2133 0.4356 0.1322 0.1940 0.3273 0.1298 0.1728
EpiMine 0.7121 0.2207 0.3243 0.7098 0.2846 0.3453 0.6267 0.2154 0.2923
- No Confidence 0.6197 0.3019 0.3845 0.4366 0.2078 0.2776 0.6029 0.2773 0.3385
Table 3. Ablation studies conducted on top 25% of article episode clusters (Section 4.3).
Terrorism Natural Disasters Politics
Ablations 5-prec 5-recall 5-F1 5-prec 5-recall 5-F1 5-prec 5-recall 5-F1
EpiMine-Top 0.2292 0.2435 0.2144 0.3817 0.2232 0.2450 0.1051 0.2233 0.1201
TF-IDF 0.0985 0.1403 0.1059 0.3284 0.1919 0.2221 0.0907 0.1908 0.0916
No DC 0.1968 0.1752 0.1707 0.2520 0.1546 0.1785 0.1126 0.2108 0.1299
Table 4. Compares the top-5 salient terms which (1) have the highest cosine-similarity (CS) and (2) highest discriminative co-occurrence (DC), with the given keyword.
Keyword CS DC
broke stormed, ransacked, dashed, occupied, rushed glass, doors, metal, building, teargas
slogans spray, placards, painted, defaced, pictures reads, wall, damage, started, portraits, spray
Table 5. Gold and detected episodes (a maximum of five are included for brevity) for the “2019 Hong Kong Legislative Protests” key event. We specify the gold/detected episode attributes for each episode cluster in the following semicolon-separated format: core entity; action; object; time; location. “Not detected” denotes that no more episodes were generated by the model.
Model Episode #1 Episode #2 Episode #3 Episode #4 Episode #5
Gold
(total 5 eps)
Activists; headed; towards the Legislative Council Complex; 1 July 2019; Hong Kong Protesters; stormed; the Legislative Council Complex; around 9:00 pm; Hong Kong; Protesters; damaged/defaced; portraits, furniture, emblem, etc.; 1 July 2019; Legislative Council Complex Police; started using; tear gas to disperse protesters; 12:05 am 2 July; around the Legislative Council complex Police; arrested; individuals in connection with the incident; between 3 July and 5 July; Hong Kong
K-means + LLM
(total 4 eps)
Protesters; storm and vandalize; Legislative Council building; July 1, 2019; Legislative Council complex in Admiralty, Hong Kong Hong Kong government; condemns; protesters storming legislative building; July 1, 2019; Hong Kong Hong Kong protesters; express; demands for freedom and democracy; July 1, 2019; Hong Kong Legislative Council Hong Kong police; adopt; more restrained tactics; July 1, 2019; Hong Kong Legislative Council Not detected
EvMine +
LLM
(total 4 eps)
Protesters; vandalize; Hong Kong legislative building; July 1, 2019; Hong Kong legislative building Protesters; occupy and vandalize; Hong Kong legislative chamber; July 1, 2019; Hong Kong legislative building Protesters; spray paint; slogans and demands; July 1, 2019; Hong Kong legislative building Protesters; deface; Hong Kong emblem; July 1, 2019; Hong Kong legislative building Not detected
Claude
(total 3 eps)
Protesters; storm; Hong Kong’s Legislative Council; July 1, 2019; Hong Kong’s Legislative Council building Police; retreat and avoid confrontation; protesters storming Hong Kong’s Legislative Council; July 1, 2019; Hong Kong’s Legislative Council building Brian Leung Kai-ping; pulls off mask and reads protesters’ demands; inside Hong Kong’s Legislative Council; July 1, 2019; Legislative Council chamber Not detected Not detected
GPT-4
(total 2 eps)
Hong Kong protesters; storm Legislative Council; government and police; July 1, 2019; Legislative Council Complex, Hong Kong Hong Kong citizens; march against extradition bill; ”Carrie Lam and Chinese government; June 2019; Various locations in Hong Kong Not detected Not detected
EpiMine (total 7 eps) protesters; broke into and occupied; Hong Kong’s legislative building; July 1, 2019; Hong Kong protesters; vandalized; the legislative building; after breaking in; Hong Kong police; fired tear gas at; protesters; after midnight on July 1; outside the legislative building Carrie Lam; condemned; the protesters’ actions; in a news conference at 4am on July 2; Hong Kong police; began making arrests of; protesters involved; in the days after; Hong Kong
5.1. Datasets
We conduct our experiments on three thematic, real-world news corpora selected from Wikipedia [^3] https://en.wikipedia.org/wiki/ over the last decade. For each theme, we manually collect approximately $10$ key events composed of multiple articles and ensure that distinct episodes exist in these articles. These articles are originally obtained from the hyperlinks of references from the Wikipage of each key event, and then filtered with constraints in time, language, and relevance. We avoid excessive filtering (allowing for articles that may contain either a subset or even none of the ground truth episodes identified) in order to preserve the challenging, real-world nature of the task.
During the collection process, we targeted selecting a diverse set of key events topics within a theme. For instance, we attempted to cover every type of “natural disaster”, including tornados, wildfires, and etc. When selecting key events, we leave out those with less than $20$ hyperlinks in the Wikipage. Table 1 summarizes the statistics for these datasets. We also construct a background news corpus of approximately 4,000 long news articles using the New York Times corpus for topic categorization (Meng et al., 2020). We provide further details on the dataset collection criteria/process, each theme, and corresponding key events in Appendices E and F.
5.2. Baseline Models
We compare against the following methods using the evaluation metrics specified in Appendix D. For the two most competitive baselines, K-means and EvMine, we also integrate EpiMine’s same exact LLM model/prompt (Appendix C) for candidate episode estimation (Section 4.3) and confidence-based episode-segment classification process (Section 4.4) to demonstrate any LLM-based improvements.
  • K-means (Likas et al., 2003): No. of ground-truth episodes is given; clusters segments based on semantic similarity of ST  (Reimers and Gurevych, 2019) embeddings.
  • EvMine (Zhang et al., 2022): Unsupervised framework for key event detection that leverages peak phrases and detects communities using event-indicative features. We extend the original document-level method to the segment level for episode detection.
  • EMiner (Jiao et al., 2023): Unsupervised event chain mining that performs atomic action clustering. For episode detection, we map its final output, a list of events, back to the original sentences from which each event was extracted, treating these sentences as segments. To retrieve more episode-associated segments, we use ST (Reimers and Gurevych, 2019) to select the $k$ most similar segments to each cluster sentence.
We also include the following full and partial ablations of EpiMine (clusters segments from all articles vs. top $\delta$articles, respectively):
  • No Confidence: A full ablation, where all input segments are classified based on the episode with max cosine similarity instead of using the confidence score from Equation 3.
  • EpiMine-Top: A partial ablation which directly outputs the intermediate episode clusters formed based on the top articles identified in Section 4.3 without inputting them into the LLM-based episode estimation step.
  • TF-IDF: A partial ablation which replaces the salience and synonym expansion step (Section 4.1) with TF-IDF.
  • No DC: A partial ablation which replaces the discriminative co-occurrence score (Equation 2) with raw pair frequency.
5.3. Overall Results & Analysis
In Table 2, EpiMine shows an average 80.8% increase in 5-precision, a 34.0% increase in 5-recall, and a 62.8% increase in 5-F1 over all baselines. Notably, despite both ’K-means’ and ’K-means + LLM’ being given the ground-truth number of episodes, they are significantly outperformed by EpiMine (both the base model and no confidence ablation). Additionally, EvMine and EMiner, originally designed for key event and atomic action levels of event granularity, fail to address the unique challenges of episode detection. We further analyze our results through extensive quantitative and qualitative studies, including a detailed case study on the “2019 Hong Kong Legislative Protest” (as shown in Figure 1), leading to the following takeaways:
  1. (1)
    The strengths of discriminative co-occurrence complement those of cosine similarity:
    Table 4 illustrates the qualitative strengths of our novel discriminative co-occurrence metric. Both cosine similarity (CS) and discriminative co-occurrence (DC) offer different, complementary strengths. CS identifies similar words that play a similar role or are synonyms within an episode (e.g., “broke”, “ransacked”), while DC identifies the key surrounding actions and objects that co-occur within the same episode (e.g., “slogans”, “wall”). This is quantitatively supported by the ablation studies in Table 3, which show a significant decrease in the quality of our top articles (partitioned into episodes) without our salience and discriminative co-occurrence measures. The politics dataset does show slight improvements in precision and F1 when discriminative co-occurrence is replaced, due to more term overlap across episodes, resulting in less distinct, sequential episodes.
  2. (2)
    LLMs require effective episode fragment clusters for quality episode estimation:
    In Table 5, we compare EpiMine with the most competitive baselines integrated with our LLM-based candidate episode estimation and classification, as well as Claude-2.1 and ChatGPT-$4$https://chat.openai.com/. Without any initial clusters as guidance, LLMs alone (Claude, GPT-4) fail to detect high-quality episodes, missing most ground-truth episodes and including irrelevant atomic actions (e.g., “Brian Leung pulls off mask”). Similarly, using low-quality clusters from the baselines results in poor performance. EvMine detects episodes that all reflect the same event, “Protesters vandalized the Legislative Chamber”. While K-means produces more distinct episodes, it does not capture the most critical, gold episodes. In contrast, EpiMine’s episodes are both distinct and meaningful.
    An interesting observation is that EpiMine’s clusters elicit the LLM to identify more meaningful temporal information. Unlike most baseline episodes, which simply have “July 1, 2019” as the time attribute, EpiMine’s episodes feature more descriptive temporal cues: “after breaking in”, “after midnight”, “in a news conference at 4 am on July 2”. Moreover, EpiMine’s “incorrect” episode #4 is a significant sub-event of the key event discussed by many articles. This strongly demonstrates the impact of EpiMine’s candidate episode clusters as input into the LLM; LLMs alone cannot perform quality episode detection.
  3. (3)
    Fragment ranking effectively identifies top articles:
    In Figure 4, we conduct a sensitivity analysis of the top articles chosen to estimate candidate episodes (Section 4.3). We compare the gold episodes contained in the set of the top $\delta$articles as we vary $\delta$. By ranking the articles based on their likelihood of containing both high-quality and numerous episodes, we find that EpiMine’s top article selection covers the vast majority of episodes by $\delta=25\%$ and more comprehensively around $\delta=45\%$. This is significant as it helps minimize both the noise and the amount of data needed to accurately detect all episodes.
  4. (4)
    Confidence leads to more conservative classification:
    The “No-Confidence” ablation in Table 2 shows that using confidence (Equation 3) improves EpiMine’s precision but reduces recall. This conservative approach results in EpiMine being cautious when assigning segments to episode clusters, excluding segments with insufficient confidence. This indicates that our method relies on both our clustering method and confidence scoring for precise episode detection. EpiMine users can determine if they prefer a confidence or no-confidence method depending on their use-case (higher precision versus higher recall). Nonetheless, both version still outperform all baselines.
Figure 4. Percentage of key event’s gold episodes captured in the $\delta$top articles chosen during the candidate episode estimation. Results averaged across themes.
6. Conclusion
In this work, we propose our novel method EpiMine for unsupervised episode detection for large-scale news events. EpiMine performs (1) episode indicative term mining, which identifies combinations of salient terms that are likely to discriminatively co-occur within an episode and not across episodes, (2) episode partitioning, which partitions each article into approximate isolated episodes based on consecutive shifts in the term co-occurrence distribution, (3) candidate episode estimation, which clusters the top partitions into candidate episodes and utilizes LLM-based reasoning to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters. EpiMine significantly outperforms all baselines on the vast majority of key events, as shown through extensive quantitative and qualitative analysis. Further work towards the temporal analysis of episodes within articles can be explored, as well as extending our work to primarily multilingual news settings with low resources.
Appendix A Ethics Statement
Based on our current methodology and results, we do not expect any significant ethical concerns, given that subtasks like episode detection within the news event extraction and analysis is a standard problem domain across data mining applications. Furthermore, having the method rely on zeor supervision helps as a barrier to any user-inputted biases. However, one minor factor to take into account is any hidden biases that exist within the large language models used as a result of any potentially biased data that they were trained on. We used these pre-trained language models for refining the fluency of the detected episode clusters and did not observe any concerning results, as it is a low-risk consideration for the domains that we studied.
Appendix B Additional Details for Episode Partitioning
We note that for determining semantic similarity between the terms of two segments, we use both (1) the average cosine similarity between all unordered pairs of terms between segment (i-1) and (i), and (2) the cosine similarity between the average of static term representations in $(i-1)$ and the average of static term representations in $(i)$. Furthermore, we filter out any non-salient segments before episode partitioning to avoid any influence of noisy segments (e.g., journalist’s analysis, summary statements, historical comparisons, and other generic noise) on the quality of our episode fragments.
Finally, following (Wang et al., 2021; Kargupta et al., 2023), we take the harmonic mean of all pairwise discriminative co-occurrence scores instead of a simple average. This allows us to prioritize the more salient and discriminative terms when determining the episode partitions. For instance, if “protesters” consistently occurs throughout the majority of episodes and thus has a low average discriminative co-occurrence, then it is not as informative for episode partitioning.
Appendix C Claude-2 Prompt & Example for Candidate Episode Estimation
Prompt. We use the following prompt for estimating fluent candidate episodes from our input episode fragment clusters. We denote $k$ as the number of episode fragment clusters outputted after clustering the top article episode fragments in Section 4.3.
Task: You are a key news event analyzer that is aiming to detect episodes (a representative subevent that reflects a critical sequence of actions performed by a subject at a certain and/or location) based on text segments from different news articles. Given the above groups of article segments, predict at least 2 and at most {k} potential episodes of the key event. Some groups may fall under the same episode. Output your answer inside the tags <answer></answer> as a JSON object where each item is also a JSON with the key "title" with the value containing the [subject, action, object, time, location] of the episode, a key "keywords" with the string value being a list of 5-10 associated keywords unique to that specific episode, and a final key "example_sentences" with a value being a list of 2-5 extracted sentences from the input segment groups. Feel free to output less than {k} episodes if you feel that any are redundant (could fit under an existing candidate episode). The title, keywords, and example sentences of a predicted episode should not be able to be placed under another different predicted episode.
Example. Below, we provide an example of EpiMine’s candidate episode estimation step (Section 4.3). Specifically, the LLM identifies the core attributes (subject, action, object, time, location) of each unique cluster, relevant keywords, and top extracted text segments given the input clusters:
title’: [’Protesters’, ’storm and vandalize’, ’Hong Kong’s Legislative Council building’, ’July 1, 2019’, ’Legislative Council building in Admiralty, Hong Kong’]
keywords’: [’vandalism, graffiti, violence, escalation, ransacking’]
example_sentences’: [’Hundreds of anti-extradition bill protesters finally broke into the legislature after many hours of attacking the public entrance and ransacked the building, including displaying the colonial Hong Kong flag in the chamber.’, ’Slogans on the wall read: ‘‘Murderous regime’’, and ‘‘There are no rioters only a tyrannical regime.’’]
Appendix D Evaluation Metrics
Following a recent work on key event detection (Zhang et al., 2022), we adapt the $k$-prec, $k$-recall, and $k$-F1 to quantitatively evaluate the episode detection results. We use these metrics to evaluate how the model output matches the ground truth episodes using the top-$k$ segments within each detected episode. Formally, suppose there are $N$ ground truth episodes $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$, each of which is a set of text segments related to its corresponding episode. =$\mathcal{E}=\{E_{1},E_{2},\dots,E_{K}\}$ are the model predicted episodes, each of which is a ranked list of segments, and $E_{j,k}$ means the top-$k$ segments within $E_{j}$. Then, the $k$-metrics are defined as follows:
k-prec $=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{\sum_{E_{j}\in\mathcal{E}}\mathds{1}(|E_{j}|\geq k)}$
k-recall $=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{N}$
k-F1 $=\frac{2\cdot\text{k-prec}\cdot\text{k-recall}}{\text{k-prec}+\text{k-recall}}$
Appendix E Key Event Corpus Retrieval
Given that our task is novel and no large-scale key event-specific news corpus is available for this task where the key events are guaranteed to contain distinguishable episodes, we briefly discuss how we collect the input corpus from online news data, even if it is out of scope for our fundamental methodology. Given our set of key events (as listed in Section 5.1), we first scrape the external reference list from their corresponding Wikipedia page and select the news articles that have been published within two months of given key event’s start date (e.g., all articles selected for “January 6 2021 Capitol Attack” would have been published between November 6-March 6). This is important as we want to prioritize the news articles which focus on describing the episodes of the key event and their corresponding aspects as opposed to primarily opinions or analyses. This allows us to motivate our task as one critical for currently evolving key events which required a more fine-grained episodic timeline. Furthermore, it is consequently unlikely for a single article to cover all of the episodes and exclusively episodes under a key event. Despite this being more challenging, it is acceptable as the goal of our task is to extract only the key event-related episodes, which must be substantiated by multiple documents in either case. We list all of our selective themes and their corresponding key events included in the dataset that we construct:
  • Terrorism and attacks: 2021 Atlanta spa shootings; 2014 Montgomery County Shootings; 2021 Indianapolis FedEx shooting; 2022 Cincinnati FBI field office attack; 2019 Jersey City shooting; 2019 Naval Air Station Pensacola shooting; 2022 Greenwood Park Mall shooting, 2018 Capital Gazette shooting; 2021 Collierville Kroger shooting; 2019 Kyoto Animation arson attack
  • Natural disasters: 2023 Tornado outbreak sequence; 2023 Hawaii Wildfires; 2021 Western Kentucky tornado; 2017 Mocoa landslide; 2010 Haiti earthquake; 2021 Henan floods; 2019 Nyonoksa radiation accident; 2022 North American winter storm; 2011 Fukushima nuclear accident
  • Political Events: 2020 Kyrgyz Revolution, 2019 Storming of the Hong Kong Legislative Council Complex; 2019 Siege of the Hong Kong Polytechnic University; 2017 Zimbabwean coup; 2018 Italian government formation; 2021 January 6 United States Capitol attack; 2018 Thai Cave Rescue Operation; 2018 Armenian Revolution; 2017 Lebanon–Saudi Arabia dispute; 2013 Tunisian political crisis
Appendix F Claude-2 Prompt for Dataset Annotation
We automatically annotate our dataset using Claude-2.1 using the prompt below (before an additional human-verification stage):
You are a news event analyzer that labels text segments of a news article with their matching event episode description. I will give you several text segments, and several episodes of a key event in tuples. We define an episode as the following: an episode is a set of thematically coherent text segments discussing a particular set of core entities performing actions for or towards an object(s) at a certain time and/or location during a real-world key event. The entities, actions, objects, time, and location can all be considered aspects of an episode.
[one-shot demonstration & format specification]
Please help classify the text segments under different episodes (the output value for each segment should be an integer key of each episode). If you think a text segment cannot be used to describe any episodes, please use "X" in the output to indicate the lack of an episode tuple number for that segment. If a text segment is very general, does not describe the key event at hand, or can be matched to multiple episodes, then please use a "M" in the output to indicate the multiple episode mapping for that segment. There should be a value assigned to each of the len(segments) segments (segment_0, ..., segment_len(segments)-1).

### v2

@@
-Unsupervised Episode Detection for Large-Scale News Events
-Abstract.
-Episodic structures are inherently interpretable and adaptable to evolving large-scale key events.
-However, state-of-the-art automatic event detection methods overlook event episodes and, therefore, struggle with these crucial characteristics.
-This paper introduces a novel task, episode detection, aimed at identifying episodes from a news corpus containing key event articles.
-An episode describes a cohesive cluster of core entities (e.g., “protesters”, “police”) performing actions at a specific time and location.
-Furthermore, an episode is a significant part of a larger group of episodes under a particular key event.
-Automatically detecting episodes is challenging because, unlike key events and atomic actions, we cannot rely on explicit mentions of times and locations to distinguish between episodes or use semantic similarity to merge inconsistent episode co-references.
-To address these challenges, we introduce EpiMine, an unsupervised episode detection framework that (1) automatically identifies the most salient, key-event-relevant terms and segments, (2) determines candidate episodes in an article based on natural episodic partitions estimated through shifts in discriminative term combinations, and (3) refines and forms final episode clusters using large language model-based reasoning on the candidate episodes.
-We construct three diverse, real-world event datasets annotated at the episode level.
-EpiMine outperforms all baselines on these datasets by an average 59.2% increase across all metrics.
-Unsupervised Episode Detection, Event Detection, Segment Classification, Large Language Models, Key Term Extraction
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Information extractionccs: Information systems Data mining
-Figure 1. Example of an event structure hierarchy. We take a key event corpus as input and aim to detect each episode node with a corresponding cluster of relevant text segments.
-1. Introduction
-Given the saturation of real-time news accessible at our fingertips, reading and processing the critical information of a key event has become an increasingly daunting challenge.
-Consequently, research on automatic textual event detection has recently attempted to integrate the manner in which humans neurologically perceive/store events into textual event detection methods.
+Synergizing Unsupervised Episode Detection with LLMs for Large-Scale News Events
+Abstract
+State-of-the-art automatic event detection struggles with interpretability and adaptability to evolving large-scale key events—unlike episodic structures, which excel in these areas.
+Often overlooked, episodes represent cohesive clusters of core entities performing actions at a specific time and location; a partially ordered sequence of episodes can represent a key event.
+This paper introduces a novel task, episode detection, which identifies episodes within a news corpus of key event articles.
+Detecting episodes poses unique challenges, as they lack explicit temporal or locational markers and cannot be merged using semantic similarity alone.
+While large language models (LLMs) can aid with these reasoning difficulties, they suffer with long contexts typical of news corpora.
+To address these challenges, we introduce EpiMine, an unsupervised framework that identifies a key event’s candidate episodes by leveraging natural episodic partitions in articles, estimated through shifts in discriminative term combinations.
+These candidate episodes are more cohesive and representative of true episodes, synergizing with LLMs to better interpret and refine them into final episodes.
+We apply EpiMine to our three diverse, real-world event datasets annotated at the episode level, where it achieves a 59.2% average gain across all metrics compared to baselines.
+Synergizing Unsupervised Episode Detection with LLMs
+for Large-Scale News Events
+Priyanka Kargupta  Yunyi Zhang  Yizhu Jiao  Siru Ouyang  Jiawei Han  Siebel School of Computing and Data Science  University of Illinois at Urbana-Champaign {pk36,yzhan238,yizhuj2,siruo2,hanj}@illinois.edu
+1 Introduction
+Given the saturation of real-time news accessible at our fingertips, reading and processing a key event’s critical information has become an increasingly daunting challenge.
+Consequently, recent work on automatic textual event detection has attempted to integrate the manner in which humans neurologically perceive/store events into textual event detection methods.
 Specifically, neuroscientists studying event representations in human memory find that events are stored in a top-to-bottom hierarchy, as demonstrated in Figure 1.
-The deeper the hierarchical event level, the more fine-grained its corresponding text granularity (Zhang et al., 2022): we consider a theme as corpus-level (all articles discussing the 2019 Hong Kong Protests), key event as document-level (an article typically discusses a full one to two day key event), episode as segment-level, and atomic action as sentence or phrase-level.
-Furthermore, neurological research (Baldassano et al., 2017; Khemlani et al., 2015) indicates that events are encoded into memory as episodic structures.
-Representing events as discrete episodes helps us piece together a coherent narrative by considering the sequence of actions, reactions, and developments over time.
-This empowers several downstream tasks (e.g., event schema generation, event prediction), which may benefit from insights into the causes and consequences of events that might be missed with a more generalized analysis.
+The deeper the hierarchical event level, the more fine-grained its corresponding text granularity Zhang et al.
+(2022): we consider a theme as corpus-level (all articles discussing the 2019 Hong Kong Protests), key event as document-level (an article typically discusses a full one to two day key event), episode as segment-level, and atomic action as sentence or phrase-level.
+Figure 1: Example event structure hierarchy. Given a key event node’s corpus, detect its episode children and their respective relevant text segments.
+Furthermore, neurological research Baldassano et al.
+(2017); Khemlani et al.
+(2015) indicates that events are encoded into memory as episodic structures.
+Representing events as discrete episodes helps us piece together a coherent and concise narrative by focusing on meaningful clusters of actions, reactions, and developments, rather than examining each in isolation or as a whole.
 Despite its strengths, existing automatic event extraction works fail to consider the episode-level.
-For example, key event detection specifically seeks to output “a set of thematically coherent documents” for each key event (Zhang et al., 2022; Liu et al., 2023).
-However, it is challenging to manually parse through a large cluster of relevant articles in order to gain an efficient and compact understanding of a given event.
-To address the lack of interpretability of such article clusters, the adjacent task of timeline summarization (Steen and Markert, 2019; Li et al., 2021b; Gholipour Ghalandari and Ifrim, 2020; Chen et al., 2023) aims to identify the dates and a compact summary for each key event.
-However, high-level timelines are typically applicable for historical themes and thus unrealistic for currently evolving key events where fine-grained timeline summarization is more suitable.
-Hence, event chain mining (Jiao et al., 2023) attempts to address this by mining a series of temporally-ordered atomic actions at the phrase-level; however, the granularity level is often too fine-grained to properly represent a large-scale key event.
-Thus, we aim to tackle the novel task of episode detection to pave the way for a more effective event representation.
+For instance, key event detection focuses on identifying “a set of thematically coherent documents” for each key event Zhang et al.
+(2022); Liu et al.
+(2023), but manually parsing large clusters of articles is inefficient and lacks interpretability.
+Timeline summarization Steen and Markert (2019); Li et al.
+(2021a); Gholipour Ghalandari and Ifrim (2020); Chen et al.
+(2023) addresses this by providing dates and compact summaries, yet it suits historical themes better than evolving key events that require finer granularity.
+Event chain mining Jiao et al.
+(2023) takes a fine-granularity approach by identifying temporally ordered atomic actions, but its phrase-level granularity is often too fine and practically redundant for large-scale events (e.g., in Figure 1, the actions all describe the same episode).
+To bridge this gap, we propose the novel task of episode detection to pave the way for a more effective event representation.
 Episode detection aims to detect episodes from a news corpus containing key event articles.
-An episode can be described as a cohesive cluster of subjects performing actions at a certain time and location, occurring as part of a larger sequence of episodes under a specific key event.
-Episode detection introduces a unique set of challenges, which we address using our novel framework, EpiMine.
-EpiMine is an unsupervised episode detection framework that automatically detects meaningful episodic events and their corresponding text segments in a large key event corpus, all without any level of human supervision or labeled training data.
-EpiMine is comprised of the following components: (1) discriminative co-occurrence detection, (2) episode partitioning, (3) candidate episode estimation, and (4) episode-segment classification.
+An episode can be described as a cohesive cluster of potentially diverse subjects performing actions at a certain time and location, occurring as part of a larger sequence of episodes under a specific key event.
+We introduce EpiMine, which detects meaningful episodic events and their corresponding text segments in a large key event corpus, all without any level of human supervision or labeled training data.
+EpiMine consists of: (1) episode indicative term mining, (2) episode partitioning, (3) LLM-enhanced episode estimation, and (4) episode-segment classification.
 Collectively, they tackle the unique challenges of episode detection, detailed below:
-Challenge 1: Journalists do not timestamp episodes.
-Key event detection partitions a thematic corpus into document-level clusters by heavily relying on explicit temporal features, like publication dates (Zhang et al., 2022).
-For example, an article primarily discusses one key event, which can then be roughly mapped to the article’s publication date.
+Challenge 1: Episodes are not timestamped.
+Key event detection partitions a thematic corpus into document-level clusters by heavily relying on explicit temporal features, like publication dates, being associated with the key event articles Zhang et al.
+(2022).
 However, this assumption fails at the episode-level, where there is no guarantee to have a distinct timestamp associated with each text segment that discusses a new episode.
-Fortunately, we can take advantage of the idea that journalists naturally partition news articles by sequentially discussing distinct episodes: {adjustwidth}1em0em Example: An article likely completes its discussion of the episode A, protesters storming the Legislaive Council, before episode B, “protesters vandalized the Legislative Chamber” (Figure 3).
-Hence, in order to partition articles into distinct episode segments, EpiMine must identify whether or not two consecutive segments are discussing the same or different episodes, which brings us to our next challenge.
+Fortunately, we can take advantage of the idea that journalists naturally partition news articles by sequentially discussing distinct episodes: {adjustwidth}0.5em0em Example: An article likely completes its discussion of the episode A, protesters storming the Legislative Council, before episode B, “protesters vandalized the Legislative Chamber” (Figure 3).
+Hence, to partition articles into distinct episode segments, EpiMine must identify whether two consecutive segments are discussing the same or different episodes— bringing us to our next challenge.
 Challenge 2: Episodes contain semantically diverse actions.
-Each episode features a set of unique atomic actions, which we can utilize for determining whether or not two segments discuss the same episode.
-However, for clustering actions, existing methods (Jiao et al., 2023) rely heavily on semantic similarity.
-This not realistic for episode-segment clustering: {adjustwidth}1em0em Example:protesters spray-painted slogans” and “they unfurled the colonial-era flag” will fall under the same episode, but are semantically different and unlikely to be clustered.
-Alternatively, we can identify salient terms that reflect the same episode (“barriers” and “shoved” are unique to Episode A; “defaced” and “walls” for Episode B), by exploiting corpus-level signals.
-Specifically, if we frequently see “defaced” mentioned together with “walls” (or their respective synonyms) and not with other terms (and vice-versa), then we consider them a discriminative co-occurrence.
-Consequently, when such co-occurrences remain consistent across text segments, this indicates the same episode being discussed.
-Conversely, if a sufficient shift in the combination of terms occurs, then this indicates that a different episode is being discussed.
+Each episode features a set of unique atomic actions, which can help determine if two segments discuss the same episode.
+However, for clustering actions, existing methods Jiao et al.
+(2023) rely heavily on semantic similarity.
+This is not realistic for episode-segment clustering: {adjustwidth}0.5em0em Example:protesters spray-painted slogans” and “they unfurled the colonial-era flag” will fall under the same episode, but are semantically different and unlikely to be clustered.
+Alternatively, we can identify notable (salient) terms unique to the same episode (episode A: “barriers” and “shoved”; episode B: “defaced” and “walls”), by exploiting corpus-level signals.
+For example, if “defaced” and “walls” are frequently mentioned together across the corpus (or their respective synonyms) and not with other terms, then they are a discriminative co-occurrence.
+When terms between two segments discriminatively co-occur, this indicates the same episode is being discussed.
+Conversely, if a sufficient shift in term combinations occurs, then a different episode is being discussed.
 Challenge 3: Articles often do not feature all episodes.
-Given the real-time nature of reporting, an article may feature only a subset or none of ground-truth episodes from a multi-day key event (e.g., focusing on a high-level analysis of the key event or mentioning other related key events).
-To minimize this noise, EpiMine seeks to identify the set of articles which maximizes the quantity and quality of potential episodes.
-It then merges any article partitions across these articles which likely discuss the same episode and employs a large language model (LLM) to provide a more fluent interpretation of the candidate episodes, accounting for the episode’s core entity, actions, object, location, and time period.
+Real-time news reporting often provides an incomplete coverage of multi-day events, with individual articles potentially omitting or partially addressing key episodes.
+Consequently, while LLMs could assist with the first two challenges, requiring multiple articles hinders their use given their long context limitations Li et al.
+(2024); Liu et al.
+(2024).
+To address this challenge, EpiMine seeks to select a minimal set of articles that maximizes both the quantity and quality of event episodes.
+It then merges any article partitions across these articles which likely discuss the same episode and synergizes with an LLM to provide a more fluent interpretation of the candidate episodes, accounting for the episode’s core entity, actions, object, location, and time period.
 This allows EpiMine to finally map the remaining non-salient article segments to these episodes, pruning any candidates which are not sufficiently supported by the remaining articles.
 We summarize our core contributions:
-  1. (1)
-    We introduce the novel task of episode detection, which takes in a key-event corpus and outputs multiple detected episodes and their related segments extracted from the corpus.
-  2. (2)
-    We propose a novel unsupervised episode detection method, EpiMine, which exploits discriminative term co-occurrences to estimate candidate episode partitions from top articles.
-It refines these candidate episodes using LLM-based reasoning to output a comprehensive and diverse set of episodes.
-  3. (3)
-    We construct three novel datasets, reflecting a diverse set of real-world themes and thirty global key events, as no large-scale key event-specific news corpus exists for this task where the key events are guaranteed to contain distinguishable episodes.
-  4. (4)
-    We compare EpiMine with five other baselines through an extensive set of experiments and case studies performed on our real-world datasets, demonstrating that it outperforms all baselines by, on average, a 59.2% increase across all metrics.
+  • Episode detection: novel task to detect episodes & their segments from a key-event corpus.
+  • EpiMine, an unsupervised episode detection method which introduces discriminative term co-occurrence and episode partitioning.
+  • Three novel datasets, reflecting a diverse set of real-world themes and thirty global key events (no key event corpus exists for this task).
+  • EpiMine outperforms all baselines by, on average, a 59.2% increase across all metrics.
 Reproducibility: We provide our dataset and source code [^1] https://github.com/pkargupta/epimine to facilitate further studies.
-2. Related Works
-2.1. Event Extraction
-Understanding events in unstructured texts is crucial, leading to substantial research in event detection and extraction (Liu et al., 2018a; Du and Cardie, 2020; Li et al., 2021a; Lu et al., 2021; Qi et al., 2022; Jiao et al., 2022).
-Existing work, such as Han et al.
+2 Related Works
+2.1 Event Extraction
+Event extraction has been widely studied, focusing on event detection Liu et al.
+(2018a); Du and Cardie (2020); Li et al.
+(2021b); Lu et al.
+(2021); Qi et al.
+(2022); Jiao et al.
+(2022), event relation extraction Han et al.
 (2019); Wang et al.
 (2020); Ahmad et al.
-(2021), have focused on event relation extraction and prediction, while salient event identification remains a significant area of study (Liu et al., 2018b; Jindal et al., 2020; Wilmot and Keller, 2021).
-Furthermore, event process understanding has garnered recent attention (Zhang et al., 2020; Chen et al., 2020).
-However, these studies typically require costly expert annotations.
-Some efforts have sought to mitigate this challenge through unsupervised approaches (Weber et al., 2018; Li et al., 2020).
-The pioneering work (Chambers and Jurafsky, 2008) introduced event chains as a novel representation of structured news knowledge.
-In the realm of biological reading comprehension, Berant et al.
-(2014) extracted events and their relationships in biological processes.
-More recently, Zhang et al.
-(2021) utilized event graphs constructed from large corpora to derive event chains for narrative understanding.
-Despite these advancements, most existing studies concentrate on phrase-level or sentence-level events (analogous to actions in this paper), neglecting the hierarchical structures of events at varying granularities within textual data.
-Additionally, these approaches often rely on human-curated event ontologies, are restricted to predefined event types, and fail to capture events from open-domain texts.
-Although key event extraction has been explored in some studies (Zhang et al., 2022; Jiao et al., 2023), the automatic detection of episodes remains under-explored and critically important.
-2.2. Timeline Summarization
-Timeline summarization (TLS) automatically identifies key dates of major events and provide concise descriptions of occurrences on those dates.
-Previous approaches to TLS have predominantly focused on extractive methods.
-For instance, Nguyen et al.
-(2014) built thematic timelines for general-domain topics by selecting and ranking events relevant to an input query.
-Martschat and Markert (2018) introduced a temporally sensitive submodularity framework, designing objective functions and constraints that model the temporal dimension inherent in timeline summarization.
-Steen and Markert (2019) were the first to propose an abstractive timeline summarization system.
-Their approach clusters sentences likely describing the same event and uses multi-sentence compression to generate summary sentences for each cluster.
-More recently, Li et al.
-(2021b) represented news articles as event-graphs, transforming the summarization task into compressing the graph to its salient sub-graph.
-Additionally, You et al.
-(2022) presented a joint learning-based heterogeneous graph attention network, integrating event detection into a unified framework to eliminate redundant sentences.
-While these methods involve key event grouping and understanding, they typically apply to historical themes, producing higher-level timelines.
-However, these approaches may not be practical for ongoing news events, where a detailed timeline summary (distinguishing between episodes) is more appropriate.
-2.3. Topic Discovery
-The goal of topic discovery is to group news articles by their themes, an area of study dating back to the early days of automated event theme identification from textual data (Allan et al., 1998).
-One approach (Zhou et al., 2015; Beykikhoshk et al., 2018) treats this task as a topic modeling issue, using LDA (Blei et al., 2003) or its enhanced versions.
-Another line of research (Sayyadi et al., 2009; Sayyadi and Raschid, 2013) constructs a graph based on keyword frequency and identifies themes from keyword clusters, mapping these themes to documents via keyword-driven feature similarities.
-However, these techniques do not consider temporal information, limiting them to discerning different thematic events.
-To address this, some studies (Fung et al., 2005; He et al., 2007) introduce tracking keyword spikes to pinpoint events.
-Specifically, (Zhao et al., 2012) uses burst detection (Kleinberg, 2004) to identify and utilize bursty terms for document representation and clustering.
-Building on this, (Ge et al., 2016) creates a burst information network with recognized bursts and their attributes for clustering.
-Finally, (Yoon et al., 2023) proposes unsupervised stream-based story discovery, which computes article embeddings their shared temporal themes.
-Overall, these methods solely focus on the key-event or atomic action level.
-Figure 2. We detail the overall framework of EpiMine.
-3. Preliminaries
-3.1. Problem Definition
-Definition 3.1 (Episode).
-An episode $E_{i}$ is one of a partially ordered sequence of subevents, $\{E_{1},\ldots,E_{i},\ldots E_{k}\}$, of a key (major) event $E$, where typically $2\geq k\geq 20$, and $E_{i}$ does not overlap with $E_{j}$ if $i\neq j$.
-Actions in the episode $E_{i}$ can be either semantically similar or diverse, but typically have relatively tight time, location, and/or thematic ( entities, actions, objects) proximity between them.
+(2021), and salient event identification Liu et al.
+(2018b); Jindal et al.
+(2020); Wilmot and Keller (2021).
+Recent work has also addressed event process understanding Zhang et al.
+(2020); Chen et al.
+(2020), though these often rely on expensive expert annotations.
+Some studies have introduced unsupervised methods to address annotation challenges Weber et al.
+(2018); Li et al.
+(2020).
+Some overlapping work exists in topic discovery, where Yoon et al.
+(2023) proposes unsupervised stream-based story discovery— computing article embeddings based on their shared temporal themes.
+Recently, large language models have demonstrated powerful general and event extraction-specific reasoning abilities Pai et al.
+(2024); Gao et al.
+(2024).
+However, traditional and LLM-driven methods either, (1) focus on phrase/sentence-level events (analogous to actions in Figure 1), or (2) require human-curated event ontologies, often overlooking interpretable, yet meaningful granularities and open-domain texts, which go beyond pre-defined event types.
+While unsupervised granular event extraction has been explored Zhang et al.
+(2022); Jiao et al.
+(2023) at the document and phrase-level, episode detection is a more interpretable granularity that remains a largely unexplored, yet vital area.
+2.2 Timeline Summarization
+Timeline summarization (TLS) identifies key dates and concise descriptions for major events.
+Early methods were extractive, focusing on ranking events for thematic timelines Nguyen et al.
+(2014) or using submodular frameworks to model temporal dimensions Martschat and Markert (2018).
+Abstractive methods later emerged, such as sentence clustering and multi-sentence compression Steen and Markert (2019).
+More recent approaches are graph-based, such as event-graph representations for salient sub-graph compression Li et al.
+(2021a) and heterogeneous GATs for redundancy reduction You et al.
+(2022).
+While they effectively summarize key events as high-level timelines, they focus on historical themes.
+Episode-level timelines for ongoing news remain underexplored.
+Figure 2: We detail the overall framework of EpiMine.
+3 Preliminaries
+3.1 Problem Definition
+Definition 1 (Episode).
+An episode $E_{i}$ is one of a partially ordered sequence of subevents, $\{E_{1},\ldots,E_{i},\ldots E_{k}\}$, of a key (major) event $E$, where typically $2\leq k\leq 20$, and $E_{i}$ does not overlap with $E_{j}$ if $i\neq j$.
+Actions in the episode $E_{i}$ can be either semantically similar or diverse, but typically have relatively tight time, location, and/or thematic (entities, actions, objects) proximity.
 EpiMine aims to extract episodes from a news corpus, where an episode occurs as a significant component of a larger group of episodes that fall under a specific key event.
-For instance, in Figure 1, without knowing Episode #1, “Protesters stormed the Legislative Council Complex”, readers would not have fully understood Episode #3, “Police dispersed protesters”.
+For instance, in Figure 1, without knowing Episode #1, “Protesters stormed the Legislative Council Complex”, readers would not fully understand Episode #3, “Police dispersed protesters”.
 Hence, episodes help us understand the overall key event and are especially useful for events that are currently evolving, where finer context is required for sufficiently understanding them.
-Definition 3.2 (Episode Detection).
-Given a corpus $\mathcal{D}$ about one key event, where each document $d\in\mathcal{D}$ is a news article, the goal of the task is to obtain the set of text segment clusters =$\mathcal{E}=\{E_{1},E_{2},\ldots,E_{K}\}$.
-Each episode cluster $E_{i}\subset\mathcal{S}=\{s_{1}^{1},s_{2}^{1},\ldots,s_{|d|}^{|\mathcal{D}|}\}$, where $\mathcal{S}$ contains all of the text segments identified in each document $d\in\mathcal{D}$, and every two clusters do not have overlapping text segments (i.e., $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
-It is important to note that the number of episodes $K$ is not known in advance and oftentimes, a news article segment may discuss either episodes of a different key event (e.g., an episode with similar aspects that occurred in a different historical key event) or multiple episodes of the current key event.
+Definition 2 (Episode Detection).
+Given a corpus $\mathcal{D}$ about one key event, where each document $d\in\mathcal{D}$ is a news article, the task is to obtain a set of text segment clusters =$\mathcal{E}=\{E_{1},E_{2},\ldots,E_{k}\}$.
+Each episode cluster $E_{i}\subset\mathcal{S}=\{s_{1}^{1},s_{2}^{1},\ldots,s_{|d|}^{|\mathcal{D}|}\}$, where $\mathcal{S}$ contains all the text segments identified in each document $d\in\mathcal{D}$, and every two clusters do not have overlapping text segments (i.e., $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
+It is important to note that $k$, the number of episodes, is not known in advance and oftentimes, a news article segment may discuss either episodes of a different key event (e.g., an episode with similar aspects that occurred in a different historical key event) or multiple episodes of the current key event.
 Nonetheless, our goal is to detect the most relevant episodes to the current key event at hand and consequently mine the most distinctive text segments for each of these (hence our constraint of $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
-3.2. Key Event News Article Pre-Processing
-Given that the expected output for the episode detection task is a cluster of text segments, we first must segment each key event news article.
-We would like to ideally preserve both the primary aspects (e.g., core entities and their actions) and peripheral aspects (e.g., reactions to a core entity’s action) relevant to that episode, which may be helpful for cross-document episode co-reference resolution.
-In order to do this, we utilize the text segmentation method, C99 (Choi, 2000).
-Furthermore, in order to assist with the cohesiveness of the segment, we employ entity co-reference resolution before performing segmentation, which assists with retaining the context across text segments (“They surrounded the legislative building […]” $\rightarrow$ “The protesters surrounded the legislative building […]”).
-In the following section, we present our core methodology given these text segments (in their raw form, without co-references resolved) and their source articles as the primary inputs.
-4. Methodology
-In order to tackle the episode detection task, we propose a novel unsupervised framework, EpiMine.
-As shown in Figure 2, EpiMine consists of the following four core components: (1) episode indicative term mining, which identifies combinations of salient terms that are likely to discriminatively co-occur within an episode and not across episodes; (2) episode partitioning, which partitions each article into approximate isolated episodes based on consecutive shifts in the term co-occurrence distribution, (3) candidate episode estimation, which clusters the top partitions into candidate episodes and utilizes LLM-based reasoning to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters.
-4.1. Episode Indicative Term Mining
-Figure 3. Natural partition between two episodes within a key event article. An episode’s discriminative terms are bolded, while salient but non-discriminative terms are underlined.
-In the absence of supervision, our goal is to identify potential candidates for episodes.
-Episodes are often described in relation to each other and usually lack specific timestamps or locations consistently mentioned within their segments.
+4 Methodology
+To tackle episode detection, we propose a novel unsupervised framework, EpiMine.
+As shown in Figure 2, EpiMine consists of the following four core components: (1) episode indicative term mining, which identifies combinations of salient terms likely to discriminatively co-occur within an episode and not across episodes; (2) episode partitioning, which partitions each article into approximate isolated episodes based on consecutive shifts in the term co-occurrence distribution, (3) LLM-enhanced candidate episode estimation, which clusters the top partitions into candidate episodes and utilizes LLM-based reasoning to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters.
+4.1 Episode Indicative Term Mining
+Figure 3: Natural partition between two episodes in a key event article. An episode’s discriminative terms are bolded; salient non-discriminative terms are underlined.
+Lacking supervision, our goal is to identify potential candidates for episodes.
+Episodes are often described in relation to each other and usually lack timestamps or locations consistently mentioned within their segments.
 For example, the phrase “police dispersed protesters” may not have a precise timestamp because it is a response to “protesters stormed the Legislative Council Complex,” and some journalists may consider the implicit ordering adequate.
-Additionally, the same episode can be described using different entities and actions, as journalists may highlight different perspectives when reporting an episode.
+Additionally, the same episode can be described using different entities and actions— journalists may report different perspectives.
 For example, both “protesters shoved against the barricades” and “the police used pepper-spray on the protesters” describe the episode “protesters stormed the Legislative Council Complex”.
-However, they are semantically different, focusing on different core entities and actions.
-Therefore, we cannot depend on a consistent subject-action-object triple or an explicit time and location associated with each episode in the articles.
-Salient Terms.
+However, they are semantically different, focused on different core entities and actions.
+Thus, we cannot depend on a consistent subject-action-object triple or an explicit time/location mapped to each episode in the article.
 To circumvent this challenge, we exploit the idea that journalists naturally partition news articles according to episodes, forming episode fragments.
@@
 Across these episode fragments, certain salient terms are featured (e.g., protesters, legislative, vandalizing, podium).
-We modify the idea of event salience from (Jiao et al., 2023) for the task of episode detection:
-Definition 4.1 (Salience).
+We adapt the idea of event salience from Jiao et al.
+(2023) specifically for the task of episode detection.
+Definition 3 (Salience).
 A term is salient if it is (1) distinct and significant to understanding a given key event, as well as (2) frequently found in a key event’s segments and infrequently in other background/general articles.
-We define the salience score of a term $w_{i}$ within segment $s$ as the following function, where $freq(w_{i})$ is the number of key event segments that $w_{i}$ is contained in, $N_{bg}$ is the number of news articles in the background corpus we construct (using general New York Times articles), and $bgf(w_{i})$ is the number of background articles that $w_{i}$ is present in.
-(1) $\textbf{Salience}(w_{i})=\left(1+\log^{2}\left(freq(w_{i})\right)\right)\log\left(\frac{N_{bg}}{bgf(w_{i})}\right)$
-Stop words and infrequent terms ($freq(w_{i})<5$) are assigned a salience score of $-1$.
-A key event’s set of salient terms $T$ is comprised of the terms with a salience score above the mean salience across the entire vocabulary.
-In the case of infrequent synonyms used by a journalist as a stylistic choice (e.g., “demonstrations”, “marches”), we expand $T$ with terms that are similar (cosine-similarity) to their static word representations (average of its contextualized word embeddings across entire key event corpus).
+Thus, we identify a set of salient terms for each segment within the corpus (salience score details in Appendix E).
 Discriminative Co-occurrence.
-Now, each text segment within the key event corpus has a corresponding set of salient terms identified.
 In Figure 3, we can see that the first episode fragment, and Episode #1 in general, features a combination of similar terms, such as “protesters”, “barrier”, and “breach”.
 Likewise, the second episode may include a combination of terms similar to “protesters”, “spray-painting”, and/or “flag”.
-Hence, we introduce the significance of co-occurrence for episode detection.
-Despite some journalists choosing to only describe the protesters spray-painting while others focus on the protesters draping the colonial-era flag, we must be able to recognize that their respective salient terms are likely to co-occur within the same episode.
+We note that despite some journalists choosing to only describe the protesters spray-painting, while others focus on the protesters draping the colonial-era flag, we must be able to recognize that their respective salient terms are likely to co-occur within the same episode.
 However, we make a novel distinction between a co-occurrence and a discriminative co-occurrence.
-A salient term $a$ (e.g., “protesters”) may often co-occur with another salient term $b$ (“spray-painting”) within an episode.
-On the other hand, if $a$ frequently co-occurs with many other terms in various episodes (“protesters broke”), $a$ and its co-occurrences are less useful for distinguishing episodes.
+Salient terms $a$ and $b$ (e.g., “protesters” and “spray-painting”) may often co-occur within an episode.
+However, if $a$ also frequently co-occurs with many other terms in various episodes (“protesters broke”), $a$ and its co-occurrences are less useful for distinguishing episodes.
 Thus, $(a,b)$ is not a discriminative co-occurrence.
-Definition 4.2 (Discriminative Co-occurrence).
+Definition 4 (Discriminative Co-occurrence).
 A pair of terms $(a,b)$ discriminatively co-occur if (1) they frequently appear together in episode $E_{i}$, and (2) neither $a$ nor $b$ appear as frequently with other terms $w$ in other episodes $E_{\notin i}$.
-We compute the discriminative occurrence $d$ between salient term pair $(a,b)$ using the following equation:
-(2) $\begin{split}\textbf{d}(a,b)=&\log\left(\frac{freq(a,b)}{\max(\bar{f_{a}},\bar{f_{b}})}\right)\times\log\left(\frac{|T|}{max(|F_{a}|,|F_{b}|)}\right),\\ \text{where }&\bar{f_{a}}=\frac{1}{|T|}\sum\limits_{\forall w_{i}\in T}freq(a,w_{i})\text{, and}\\ &F_{a}=\{freq(a,w_{i})>1\text{ }\forall\text{ }w_{i}\in T\}\end{split}$
+We compute the discriminative occurrence $d$ between salient term pair $(a,b)$ using the following:
+$\begin{split}\textbf{d}(a,b)=&\log\left(\frac{freq(a,b)}{\max(\bar{f_{a}},\bar{f_{b}})}\right)\times\log\left(\frac{|T|}{max(|F_{a}|,|F_{b}|)}\right),\\ \text{where }&\bar{f_{a}}=\frac{1}{|T|}\sum\limits_{\forall w_{i}\in T}freq(a,w_{i})\text{, and}\\ &F_{a}=\{freq(a,w_{i})>1\text{ }\forall\text{ }w_{i}\in T\}\end{split}$ (1)
 The first $\log$ term ensures that the pair’s co-occurrence ($freq(a,b)$) is statistically significant ($\geq$ the max of $a$ and $b$’s mean vocabulary-wide co-occurrence respectively).
-The second $\log$ term ensures that the salient term pair is a discriminative matching, penalizing pairings where either $a$ or $b$ co-occurs with a large proportion of the salient term set $T$).
+The second $\log$ term ensures the pair is a discriminative match, penalizing cases where $a$ or $b$ frequently co-occurs with a large portion of the salient term set $T$.
 For example, co-occurrences with “protesters” are not discriminative because “protesters” is a core entity in all episodes and thus frequently co-occurs with many terms in $T$.
@@
 By leveraging multiple articles in a large key event corpus, we have sufficient statistical support to ensure our output reflects the average realistic reporting of the key event and its episodes.
-4.2. Episode Partitioning
+4.2 Episode Partitioning
 With the ability to identify discriminative co-occurrences, we can use a key transitive property to resolve episode co-references within and across articles, where not all combinations of an episode’s discriminative terms explicitly co-occur:
@@
 To illustrate this, we have the following text segment excerpts of a news article (the salient and discriminative terms are italicized):
-  1. (1)
-    Protesters defaced the emblem of Hong Kong, spray-painted slogans, and unfurled the colonial-era flag.
-  2. (2)
-    The portrait of LegCo president Andrew Leung was defaced.
-  3. (3)
+  1. 1.
+    Protesters defaced the Hong Kong emblem, spray-painted slogans, and unfurled the flag.
+  2. 2.
+    The portrait of LegCo president was defaced.
+  3. 3.
     A slogan on the wall reads: “The government forced us to revolt”.
-  4. (4)
+  4. 4.
     Police said at least 13 people had been arrested on suspicion of involvement in the pro-democracy protest.
@@
 If not (e.g., (slogans, arrested) for segments 3-4), this indicates that a different episode is being discussed, and we partition them into two episode fragments.
-Further implementation details are provided in Appendix B.
-4.3. Candidate Episode Estimation
-After partitioning each article into episode fragments, we can identify the set of articles that maximizes the quantity and quality of potential episodes.
-This step is crucial as it enables EpiMine to minimize noise from articles that may discuss only a subset or none of the ground-truth episodes (e.g., focusing on high-level analyses or other relevant key events).
-Simultaneously, it improves our method’s efficiency by not relying on the entire key event corpus.
-We compute the per-article rank by multiplying two metrics:
-  1. (1)
+Further implementation details are provided in Appendix G.
+4.3 LLM-Enhanced Episode Estimation
+LLMs demonstrate strong event-specific reasoning at the phrase or sentence level (Pai et al., 2024; Gao et al., 2024), but they struggle with understanding long contexts (Li et al., 2024; Liu et al., 2024).
+This limitation hampers their ability to process all episode fragments for detecting episodes.
+Additionally, noisy retrieval significantly affects reasoning performance (Shen et al., 2024).
+To address these challenges, we propose a synergistic approach that enhances in-context episode reasoning by reducing the number of required fragments while improving their cohesiveness and quality.
+We first identify the set of articles that maximizes the quantity and quality of potential episodes, where each article is ranked by multiplying two metrics:
+  1. 1.
     Quality of episode fragments: A top article should primarily consist of episode fragments containing salient terms that discriminatively co-occur.
 This reduces the rank of general fragments which summarize/analyze the event.
-We use the average of each episode fragment’s mean inner-discriminative co-occurrence (across all pairwise combinations of a fragment’s salient terms).
-  2. (2)
-    Quantity of episode fragments: A top article should ideally contain the highest number of ground-truth episodes. Therefore, we take the $\log$ of the number of episode fragments in the article.
-After ranking all articles, we select the top $\delta$to serve as the basis for candidate episode estimation.
-To estimate the candidate episodes, we must first resolve potential co-references to the same episode across these top articles.
-We apply agglomerative clustering (Murtagh and Contreras, 2012) to the top episode fragments using a pre-computed distance matrix.
+We average each episode fragment’s mean inner-discriminative co-occurrence (across all pairwise combinations of its salient terms).
+  2. 2.
+    Quantity of episode fragments: A top article should ideally contain all ground-truth episodes. Therefore, we take the $\log$ of the number of episode fragments in the article.
+After ranking all articles, we select the top $\delta$and resolve potential co-references to the same episode across these top articles.
+We apply agglomerative clustering Murtagh and Contreras (2012) to the top episode fragments using a pre-computed distance matrix.
 The distance between two fragments (inversed) is calculated using the same discriminative and static semantic similarity score used in Section 4.2).
-We then prune clusters that do not contain a statistically significant number of episode fragments.
-Finally, we use an LLM to resolve (1) the lack of time and location stamps in an episode fragment and (2) semantic inconsistency between the fragments within a cluster.
-The LLM leverages its reasoning capabilities to estimate the candidate episodes by identifying the core attributes of each unique cluster (entities, actions, objects, location, and time).
-It outputs the episode attributes, relevant keywords (assists with extraction), and top extracted text segments based on the clusters (prompt & example in Appendix C).
-4.4. Episode-Segment Classification
+Clusters with a statistically insignificant number of episode fragments are pruned.
+Finally, we provide episode fragment clusters as a more interpretable context for the LLM to resolve two challenges: (1) missing time and location stamps in fragments, and (2) semantic inconsistencies within clusters.
+The LLM summarizes each cluster by identifying its core attributes— entities, actions, objects, location, and time.
+It then outputs the episode attributes, relevant keywords for extraction, and the top extracted text segments (prompt & example in Appendix H).
+4.4 Episode-Segment Classification
 With these core summaries of the episode clusters, we obtain a generalized description of each candidate episode.
-This allows us to compute a simple episode representation for assigning an episode and confidence score to each input segment.
-Episode & Segment Representations.
-We compute a candidate episode representation by encoding both its LLM core attributes and extracted segments using SentenceTransformers (ST) (Reimers and Gurevych, 2019).
-Following extremely weakly supervised text classification works, such as (Wang et al., 2021; Kargupta et al., 2023), we take the harmonic mean of these representations, as the latter extracted segments are likely not as significant as the earlier extractions and core attributes.
-We similarly encode all input segments with the same ST model.
-Episode-Segment Confidence Estimation
-Following recent work on extremely weakly supervised text classification (Kargupta et al., 2023), we compute a confidence score specific to the episode-segment classification task.
-Simply mapping a text segment to its top episode (based on highest cosine similarity) overlooks the presence of episode-irrelevant text segments and fails to address segments discussing multiple episodes (e.g., a journalist’s key event summary at the beginning of an article).
-We aim to avoid assigning such segments to any cluster, ensuring no overlap between episode clusters, as discussed in Section 3.1.
-Thus, we compute segment $s_{i}$’s cosine similarity to its top two episodes ($e_{i}^{0}$ and $e_{i}^{1}$).
-A larger gap ($e_{i}^{0}-e_{i}^{1}$) between the two indicates higher confidence in mapping $s_{i}$ to its top episode.
-We normalize each gap by the sum of all segment episode gaps within the entire corpus, ensuring the confidence is relative to the key event:
-(3) $s_{i,\text{confidence}}=\frac{e_{i}^{0}-e_{i}^{1}}{\sum_{l=1}^{|\mathcal{S}|}(e^{0}_{l}-e^{1}_{l})}$
-After computing each segment’s confidence in its top episode, we assign only those segments with a statistically significant confidence level to their respective episode clusters $E$.
-We then prune candidate episode clusters with $\leq 0$ assigned segments, resulting in our final detected episodes and their clusters, $\mathcal{E}$.
-5. Experimental Setup
+For each candidate, we encode its LLM-based core attributes and extracted segments to compute a simple episode representation.
+Specifically, following extremely weakly supervised text classification works Wang et al.
+(2021); Kargupta et al.
+(2023), we take the harmonic mean of these representations—as the latter extracted segments are likely not as significant as the earlier extractions and core attributes.
+We similarly encode all input segments with the same encoder.
+We use these to assign an episode and confidence score to each encoded input segment.
+Episode-Segment Confidence Estimation.
+Directly mapping a text segment to its top episode based on cosine similarity risks misclassifying episode-irrelevant segments or those discussing multiple episodes (e.g., a journalist’s summary).
+To avoid classifying such segments and ensure non-overlapping episode clusters (as discussed in Section 3.1), we must determine the confidence of a segment discussing a single episode.
+We compute segment $s_{i}$’s cosine similarity to its top two episodes ($e_{i}^{0}$ and $e_{i}^{1}$).
+A larger gap ($e_{i}^{0}-e_{i}^{1}$) reflects greater confidence in classifying $s_{i}$ to $e_{i}^{0}$.
+Each gap is normalized by the sum of all segment-episode gaps across the corpus, ensuring confidence is relative to the key event:
+$s_{i,\text{confidence}}=\frac{e_{i}^{0}-e_{i}^{1}}{\sum_{l=1}^{|\mathcal{S}|}(e^{0}_{l}-e^{1}_{l})}$ (2)
+Segments with statistically significant confidence in their top episode are assigned to their respective episode clusters $E_{i}$.
+Episodes with no assigned segments are pruned, yielding the final detected episodes and clusters, $\mathcal{E}$.
+5 Experiments
 For implementing EpiMine, we use the following hyperparameters across all datasets: $\delta=25\%$, $sim\_thresh=0.75$.
+We also use Claude-2.1 as our base LLM ([Uncaptioned image]).
 All other hyperparameters are set to their respective default values.
-To determine statistical significance, we check for whether or not a value is $\geq\mu-\sigma$.
-For our word representations, we use bert-base-uncased.
-For our sentence representations, we use all-mpnet-base-v2.
-We choose to use Claude-2.1222claude.ai/ for fluent candidate episode estimation due to its strong structured JSON/XML input and output formatting abilities.
-However, this proprietary model can be replaced with any open-source model as EpiMine is model-agnostic.
-We use only one NVIDIA GeForce GTX 1080 for all experiments.
-Table 1. Statistics of our collected datasets. The numbers are averaged per key event.
+We provide all experimental settings in Appendix A.
+Table 1: Statistics of our collected datasets. The numbers are averaged per key event.
 Theme # docs # episodes # segments
@@
 Political Events 70.2 7.5 667.7
-Table 2.
-Results averaged across each theme, including the mean # of episodes EpiMine identifies per theme.
+Table 2: Results averaged across each theme, including the mean # of episodes that EpiMine identifies per theme (in parenthesis).
 Results are computed on each key event corpus using the top-5 documents for each detected episode.
 Due to variance in LLM generation, we run it 10 times and report the average of each measure.
-Bold values denote the top method; $\dagger$ denotes the second-best method.
-Terrorism (avg. # of eps = 5.36) Natural Disasters (avg. # of eps = 7.4) Politics (avg. # of eps = 7.5)
+We scale each value by 100.
+Bold values denote the top method; second-best method is underlined.
+Terrorism (5.36 eps) Natural Disasters (7.4 eps) Politics (7.5 eps)
 Methods 5-prec 5-recall 5-F1 5-prec 5-recall 5-F1 5-prec 5-recall 5-F1
-EMiner 0.0864 0.0025 0.0048 0.1037 0.0019 0.0037 0.0866 0.0016 0.0032
-K-means 0.2123 0.2123 0.2123 0.2785 0.2847 0.2814 0.1604 0.1604 0.1604
-K-means + LLM 0.2858 0.1404 0.1826 0.3740 0.1658 0.22 0.2718 0.1736 0.1825
-EvMine 0.2303 0.1502 0.1745 0.2815 0.0802 0.1225 0.0536 0.04 0.0458
-EvMine + LLM 0.3788 0.1570 0.2133 0.4356 0.1322 0.1940 0.3273 0.1298 0.1728
-EpiMine 0.7121 0.2207 0.3243 0.7098 0.2846 0.3453 0.6267 0.2154 0.2923
-- No Confidence 0.6197 0.3019 0.3845 0.4366 0.2078 0.2776 0.6029 0.2773 0.3385
-Table 3. Ablation studies conducted on top 25% of article episode clusters (Section 4.3).
+EMiner 8.64 0.25 0.48 10.37 0.19 0.37 8.66 0.16 0.32
+K-means 21.23 21.23 21.23 27.85 28.47 28.14 16.04 16.04 16.04
+K-means + [Uncaptioned image] 28.58 14.04 18.26 37.40 16.58 22.00 27.18 17.36 18.25
+EvMine 23.03 15.02 17.45 28.15 8.02 12.25 5.36 4.00 4.58
+EvMine + [Uncaptioned image] 37.88 15.70 21.33 43.56 13.22 19.40 32.73 12.98 17.28
+EpiMine ([Uncaptioned image]) 71.21 22.07 32.43 70.98 28.46 34.53 62.67 21.54 29.23
+- No Confidence ([Uncaptioned image]) 61.97 30.19 38.45 43.66 20.78 27.76 60.29 27.73 24.77
+- No LLM 37.73 21.62 24.77 37.19 14.78 17.52 30.64 23.51 19.06
+5.1 Datasets
+We conduct our experiments on three novel thematic, real-world news corpora selected from Wikipedia [^2] https://en.wikipedia.org/wiki/ over the last decade.
+For each theme, we manually collect approximately $10$ key events composed of multiple articles and ensure that distinct episodes exist in each:
+  • Terrorism and attacks: 2021 Atlanta spa shootings; 2014 Montgomery County Shootings; 2021 Indianapolis FedEx shooting; 2022 Cincinnati FBI field office attack; 2019 Jersey City shooting; 2019 Naval Air Station Pensacola shooting; 2022 Greenwood Park Mall shooting, 2018 Capital Gazette shooting; 2021 Collierville Kroger shooting; 2019 Kyoto Animation arson attack
+  • Natural disasters: 2023 Tornado outbreak sequence; 2023 Hawaii Wildfires; 2021 Western Kentucky tornado; 2017 Mocoa landslide; 2010 Haiti earthquake; 2021 Henan floods; 2019 Nyonoksa radiation accident; 2022 NA winter storm; 2011 Fukushima nuclear accident
+  • Political events: 2020 Kyrgyz Revolution, 2019 Storming of the Hong Kong Legislative Council Complex; 2019 Siege of the Hong Kong Polytechnic University; 2017 Zimbabwean coup; 2018 Italian government formation; 2021 January 6 U.S.
+Capitol attack; 2018 Thai Cave Rescue Operation; 2018 Armenian Revolution; 2017 Lebanon–Saudi Arabia dispute; 2013 Tunisian political crisis
+The articles are obtained from the Wikipage references of each key event— filtered with constraints in time, language, and relevance.
+Furthermore, each article is segmented to match our setting (Appendix F), and each segment is automatically annotated (Appendix I) with either its corresponding episode ID or with a multiple/no episode tag (‘M’ or ‘X’).
+Further details on the criteria/process, each theme, and corresponding key events are in Appendices D and I.
+We also conduct a human-automatic agreement analysis for segment-episode annotation, which shows substantial agreement with good reliability (Appendix D).
+5.2 Baselines
+We compare against the following methods using the evaluation metrics specified in Appendix C: (1) K-means Likas et al.
+(2003): given the # of ground-truth episodes, it clusters segments using ST  Reimers and Gurevych (2019) embeddings; (2) EvMine Zhang et al.
+(2022): a document-level unsupervised key event detection method adapted to segment level for episode detection; (3) EMiner Jiao et al.
+(2023): unsupervised event chain miner that clusters atomic actions, adapted to episodes; (4) No Confidence: an ablation that uses max cosine-similarity instead of confidence from Equation 2; (5) No LLM: an ablation that uses estimated episode clusters from Section 4.3 to compute our episode representations directly.
+We also integrate [Uncaptioned image] into K-means and EvMine using our same prompt (Appendix H).
+All baseline and ablation details are in Appendix B.
+Table 3: Ablation studies conducted on top 25% of article episode clusters (Section 4.3).
 Terrorism Natural Disasters Politics
@@
 No DC 0.1968 0.1752 0.1707 0.2520 0.1546 0.1785 0.1126 0.2108 0.1299
-Table 4. Compares the top-5 salient terms which (1) have the highest cosine-similarity (CS) and (2) highest discriminative co-occurrence (DC), with the given keyword.
+Table 4: Compares top-5 salient terms which (1) have the highest cosine-sim (CS) and (2) discriminative co-occurrence (DC), with the given keyword.
 Keyword CS DC
@@
 slogans spray, placards, painted, defaced, pictures reads, wall, damage, started, portraits, spray
-Table 5.
-Gold and detected episodes (a maximum of five are included for brevity) for the “2019 Hong Kong Legislative Protests” key event.
+Table 5: Gold and detected episodes (a maximum of five are included for brevity) for the “2019 Hong Kong Legislative Protests” key event.
 We specify the gold/detected episode attributes for each episode cluster in the following semicolon-separated format: core entity; action; object; time; location.
 “Not detected” denotes that no more episodes were generated by the model.
+We note the number of detected episodes beside the model name.
+We also color-code attributes which clearly align to a specific episode.
 Model Episode #1 Episode #2 Episode #3 Episode #4 Episode #5
 Gold
-(total 5 eps)
+(5 eps)
 Activists; headed; towards the Legislative Council Complex; 1 July 2019; Hong Kong Protesters; stormed; the Legislative Council Complex; around 9:00 pm; Hong Kong; Protesters; damaged/defaced; portraits, furniture, emblem, etc.; 1 July 2019; Legislative Council Complex Police; started using; tear gas to disperse protesters; 12:05 am 2 July; around the Legislative Council complex Police; arrested; individuals in connection with the incident; between 3 July and 5 July; Hong Kong
-K-means + LLM
-(total 4 eps)
+K-means + [Uncaptioned image]
+(4 eps)
 Protesters; storm and vandalize; Legislative Council building; July 1, 2019; Legislative Council complex in Admiralty, Hong Kong Hong Kong government; condemns; protesters storming legislative building; July 1, 2019; Hong Kong Hong Kong protesters; express; demands for freedom and democracy; July 1, 2019; Hong Kong Legislative Council Hong Kong police; adopt; more restrained tactics; July 1, 2019; Hong Kong Legislative Council Not detected
-EvMine +
-LLM
-(total 4 eps)
+EvMine + [Uncaptioned image]
+(4 eps)
 Protesters; vandalize; Hong Kong legislative building; July 1, 2019; Hong Kong legislative building Protesters; occupy and vandalize; Hong Kong legislative chamber; July 1, 2019; Hong Kong legislative building Protesters; spray paint; slogans and demands; July 1, 2019; Hong Kong legislative building Protesters; deface; Hong Kong emblem; July 1, 2019; Hong Kong legislative building Not detected
-Claude
-(total 3 eps)
+Claude ([Uncaptioned image])
+(3 eps)
 Protesters; storm; Hong Kong’s Legislative Council; July 1, 2019; Hong Kong’s Legislative Council building Police; retreat and avoid confrontation; protesters storming Hong Kong’s Legislative Council; July 1, 2019; Hong Kong’s Legislative Council building Brian Leung Kai-ping; pulls off mask and reads protesters’ demands; inside Hong Kong’s Legislative Council; July 1, 2019; Legislative Council chamber Not detected Not detected
 GPT-4
-(total 2 eps)
-Hong Kong protesters; storm Legislative Council; government and police; July 1, 2019; Legislative Council Complex, Hong Kong Hong Kong citizens; march against extradition bill; ”Carrie Lam and Chinese government; June 2019; Various locations in Hong Kong Not detected Not detected
-EpiMine (total 7 eps) protesters; broke into and occupied; Hong Kong’s legislative building; July 1, 2019; Hong Kong protesters; vandalized; the legislative building; after breaking in; Hong Kong police; fired tear gas at; protesters; after midnight on July 1; outside the legislative building Carrie Lam; condemned; the protesters’ actions; in a news conference at 4am on July 2; Hong Kong police; began making arrests of; protesters involved; in the days after; Hong Kong
-5.1. Datasets
-We conduct our experiments on three thematic, real-world news corpora selected from Wikipedia [^3] https://en.wikipedia.org/wiki/ over the last decade.
-For each theme, we manually collect approximately $10$ key events composed of multiple articles and ensure that distinct episodes exist in these articles.
-These articles are originally obtained from the hyperlinks of references from the Wikipage of each key event, and then filtered with constraints in time, language, and relevance.
-We avoid excessive filtering (allowing for articles that may contain either a subset or even none of the ground truth episodes identified) in order to preserve the challenging, real-world nature of the task.
-During the collection process, we targeted selecting a diverse set of key events topics within a theme.
-For instance, we attempted to cover every type of “natural disaster”, including tornados, wildfires, and etc.
-When selecting key events, we leave out those with less than $20$ hyperlinks in the Wikipage.
-Table 1 summarizes the statistics for these datasets.
-We also construct a background news corpus of approximately 4,000 long news articles using the New York Times corpus for topic categorization (Meng et al., 2020).
-We provide further details on the dataset collection criteria/process, each theme, and corresponding key events in Appendices E and F.
-5.2. Baseline Models
-We compare against the following methods using the evaluation metrics specified in Appendix D.
-For the two most competitive baselines, K-means and EvMine, we also integrate EpiMine’s same exact LLM model/prompt (Appendix C) for candidate episode estimation (Section 4.3) and confidence-based episode-segment classification process (Section 4.4) to demonstrate any LLM-based improvements.
-  • K-means (Likas et al., 2003): No. of ground-truth episodes is given; clusters segments based on semantic similarity of ST  (Reimers and Gurevych, 2019) embeddings.
-  • EvMine (Zhang et al., 2022): Unsupervised framework for key event detection that leverages peak phrases and detects communities using event-indicative features.
-We extend the original document-level method to the segment level for episode detection.
-  • EMiner (Jiao et al., 2023): Unsupervised event chain mining that performs atomic action clustering.
-For episode detection, we map its final output, a list of events, back to the original sentences from which each event was extracted, treating these sentences as segments.
-To retrieve more episode-associated segments, we use ST (Reimers and Gurevych, 2019) to select the $k$ most similar segments to each cluster sentence.
-We also include the following full and partial ablations of EpiMine (clusters segments from all articles vs. top $\delta$articles, respectively):
-  • No Confidence: A full ablation, where all input segments are classified based on the episode with max cosine similarity instead of using the confidence score from Equation 3.
-  • EpiMine-Top: A partial ablation which directly outputs the intermediate episode clusters formed based on the top articles identified in Section 4.3 without inputting them into the LLM-based episode estimation step.
-  • TF-IDF: A partial ablation which replaces the salience and synonym expansion step (Section 4.1) with TF-IDF.
-  • No DC: A partial ablation which replaces the discriminative co-occurrence score (Equation 2) with raw pair frequency.
-5.3. Overall Results & Analysis
+(2 eps)
+Hong Kong protesters; storm Legislative Council; government and police; July 1, 2019; Legislative Council Complex, Hong Kong Hong Kong citizens; march against extradition bill; "Carrie Lam and Chinese government; June 2019; Various locations in Hong Kong Not detected Not detected
+EpiMine w/ [Uncaptioned image]
+(7 eps)
+Protesters; broke into and occupied; Hong Kong’s legislative building; July 1, 2019; Hong Kong Protesters; vandalized; the legislative building; after breaking in; Hong Kong Police; fired tear gas at; protesters; after midnight on July 1; outside the legislative building Carrie Lam; condemned; the protesters’ actions; in a news conference at 4am on July 2; Hong Kong Police; began making arrests of; protesters involved; in the days after; Hong Kong
+5.3 Overall Results & Analysis
 In Table 2, EpiMine shows an average 80.8% increase in 5-precision, a 34.0% increase in 5-recall, and a 62.8% increase in 5-F1 over all baselines.
-Notably, despite both ’K-means’ and ’K-means + LLM’ being given the ground-truth number of episodes, they are significantly outperformed by EpiMine (both the base model and no confidence ablation).
+Notably, despite both K-means and K-means + [Uncaptioned image] being given the ground-truth number of episodes, they are significantly outperformed by EpiMine (both the base model and no confidence ablation).
 Additionally, EvMine and EMiner, originally designed for key event and atomic action levels of event granularity, fail to address the unique challenges of episode detection.
 We further analyze our results through extensive quantitative and qualitative studies, including a detailed case study on the “2019 Hong Kong Legislative Protest” (as shown in Figure 1), leading to the following takeaways:
-  1. (1)
-    The strengths of discriminative co-occurrence complement those of cosine similarity:
+1.
+LLMs require effective episode fragment clusters for synergistic episode estimation.
+As shown in Table 5, LLMs without any initial clusters as guidance ([Uncaptioned image] , GPT-4333https://chat.openai.com/) fail to detect high-quality episodes, miss most ground-truth episodes, and include irrelevant atomic actions (e.g., “Brian Leung pulls off mask”).
+Similarly, using low-quality baseline clusters results in poor performance.
+EvMine detects episodes that all reflect the same event, “Protesters vandalized the Legislative Chamber”.
+While K-means produces more distinct episodes, it does not capture the most critical, gold episodes.
+In contrast, EpiMine’s episodes are both distinct and meaningful, attributed to its cluster quality.
+This is quantitatively confirmed by EpiMine-No LLM’s competitive performance: using only EpiMine’s episode fragment clusters to compute episode representations—without any LLM summarization—still yields significantly better performance than all baselines (without or without LLM integration) on the Terrorism and Politics datasets, and remains highly competitive on Natural Disasters.
+This indicates the high quality of our fragment ranking and clustering.
+Figure 4: Percentage of key event’s gold episodes captured in the $\delta$top articles chosen during the candidate episode estimation. Results averaged across themes.
+EpiMine’s clusters also elicit the LLM to identify more meaningful temporal information.
+Unlike most baseline episodes which have “July 1, 2019” as the time attribute, EpiMine’s episodes feature more descriptive temporal cues: “after breaking in”, “after midnight”, “in a news conference at 4 am on July 2”.
+Moreover, EpiMine’s “incorrect” episode #4 is a significant sub-event of the key event discussed by many articles.
+This strongly demonstrates the impact of EpiMine’s candidate episode clusters as input into the LLM; LLMs alone cannot perform quality episode detection.
+2.
+The strengths of discriminative co-occurrence complement those of cosine similarity.
 Table 4 illustrates the qualitative strengths of our novel discriminative co-occurrence metric.
@@
 CS identifies similar words that play a similar role or are synonyms within an episode (e.g., “broke”, “ransacked”), while DC identifies the key surrounding actions and objects that co-occur within the same episode (e.g., “slogans”, “wall”).
-This is quantitatively supported by the ablation studies in Table 3, which show a significant decrease in the quality of our top articles (partitioned into episodes) without our salience and discriminative co-occurrence measures.
-The politics dataset does show slight improvements in precision and F1 when discriminative co-occurrence is replaced, due to more term overlap across episodes, resulting in less distinct, sequential episodes.
-  2. (2)
-    LLMs require effective episode fragment clusters for quality episode estimation:
-    In Table 5, we compare EpiMine with the most competitive baselines integrated with our LLM-based candidate episode estimation and classification, as well as Claude-2.1 and ChatGPT-$4$https://chat.openai.com/.
-Without any initial clusters as guidance, LLMs alone (Claude, GPT-4) fail to detect high-quality episodes, missing most ground-truth episodes and including irrelevant atomic actions (e.g., “Brian Leung pulls off mask”).
-Similarly, using low-quality clusters from the baselines results in poor performance.
-EvMine detects episodes that all reflect the same event, “Protesters vandalized the Legislative Chamber”.
-While K-means produces more distinct episodes, it does not capture the most critical, gold episodes.
-In contrast, EpiMine’s episodes are both distinct and meaningful.
-    An interesting observation is that EpiMine’s clusters elicit the LLM to identify more meaningful temporal information.
-Unlike most baseline episodes, which simply have “July 1, 2019” as the time attribute, EpiMine’s episodes feature more descriptive temporal cues: “after breaking in”, “after midnight”, “in a news conference at 4 am on July 2”.
-Moreover, EpiMine’s “incorrect” episode #4 is a significant sub-event of the key event discussed by many articles.
-This strongly demonstrates the impact of EpiMine’s candidate episode clusters as input into the LLM; LLMs alone cannot perform quality episode detection.
-  3. (3)
-    Fragment ranking effectively identifies top articles:
+This is quantitatively supported by TF-IDF and No DC in Table 3, which show a significant decrease in the quality of our top episode clusters without our salience and discriminative co-occurrence measures.
+We note that the politics dataset does show slight improvements in precision and F1 when discriminative co-occurrence is replaced, likely due to more term overlap across episodes (less distinct, sequential episodes).
+3.
+Fragment ranking identifies top articles.
 In Figure 4, we conduct a sensitivity analysis of the top articles chosen to estimate candidate episodes (Section 4.3).
@@
 This is significant as it helps minimize both the noise and the amount of data needed to accurately detect all episodes.
-  4. (4)
-    Confidence leads to more conservative classification:
-    The “No-Confidence” ablation in Table 2 shows that using confidence (Equation 3) improves EpiMine’s precision but reduces recall.
-This conservative approach results in EpiMine being cautious when assigning segments to episode clusters, excluding segments with insufficient confidence.
-This indicates that our method relies on both our clustering method and confidence scoring for precise episode detection.
-EpiMine users can determine if they prefer a confidence or no-confidence method depending on their use-case (higher precision versus higher recall).
-Nonetheless, both version still outperform all baselines.
-Figure 4. Percentage of key event’s gold episodes captured in the $\delta$top articles chosen during the candidate episode estimation. Results averaged across themes.
-6. Conclusion
-In this work, we propose our novel method EpiMine for unsupervised episode detection for large-scale news events.
-EpiMine performs (1) episode indicative term mining, which identifies combinations of salient terms that are likely to discriminatively co-occur within an episode and not across episodes, (2) episode partitioning, which partitions each article into approximate isolated episodes based on consecutive shifts in the term co-occurrence distribution, (3) candidate episode estimation, which clusters the top partitions into candidate episodes and utilizes LLM-based reasoning to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters.
+This is further supported by the EpiMine-Top ablation (Table 3), which quantitatively shows that the top 25% of the ranked fragments alone have competitive or higher performance than the baselines.
+The Role of Confidence.
+The confidence metric influences EpiMine toward more conservative episode-segment classification by pruning segments with statistically insignificant confidence scores ($\leq\mu-\sigma$), which are unlikely to map to a single episode.
+Reducing this threshold increases the number of mapped segments, potentially improving recall by converting false negatives into true positives, but at the cost of reduced precision due to an increase in false positives.
+This trade-off is quantitatively demonstrated in the “No-Confidence” ablation (Table 2), where omitting confidence leads to occasional gains in recall accompanied by declines in precision.
+Nevertheless, both configurations—with and without confidence—are significantly better than all baselines, allowing users to determine if including confidence aligns with their use case.
+6 Conclusion
+In this work, we proposed EpiMine, a novel, unsupervised episode detection method for large-scale news events.
+EpiMine performs (1) episode indicative term mining— identifying combinations of salient terms that are likely to discriminatively co-occur within an episode and not across episodes, (2) episode partitioning, which partitions each article into approximate isolated episodes, (3) LLM-enhanced episode estimation, which clusters the top partitions into candidate episodes and synergizes with an LLM to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters.
 EpiMine significantly outperforms all baselines on the vast majority of key events, as shown through extensive quantitative and qualitative analysis.
+7 Limitations & Future Work
+While EpiMine serves as an intuitive, unsupervised framework which demonstrates a more interpretable granularity for event analysis (episodes), it contains a few limitations that form the foundation for future, impactful research areas.
+We note that the key event theme has an impact on EpiMine’s performance.
+Specifically, natural disaster episodes are typically sequential and semantically distinct: disaster begins $\rightarrow$ warning $\rightarrow$ evacuation $\rightarrow$ damage/deaths $\rightarrow$ relief.
+As K-means is uniquely given $k$, the number of episodes, and relies on semantic similarity, it performs well with distinct episodes.
+However, we still see that its reliance on surface-level semantics leads to lower precision.
+Additionally, in the ablation studies shown in Table 3, the politics dataset does show slight improvements in precision and F1 when discriminative co-occurrence is replaced, due to more term overlap across episodes, resulting in less distinct, sequential episodes.
 Further work towards the temporal analysis of episodes within articles can be explored, as well as extending our work to primarily multilingual news settings with low resources.
-Appendix A Ethics Statement
+8 Ethics Statement
 Based on our current methodology and results, we do not expect any significant ethical concerns, given that subtasks like episode detection within the news event extraction and analysis is a standard problem domain across data mining applications.
-Furthermore, having the method rely on zeor supervision helps as a barrier to any user-inputted biases.
+Furthermore, having the method rely on zero supervision helps as a barrier to any user-inputted biases.
 However, one minor factor to take into account is any hidden biases that exist within the large language models used as a result of any potentially biased data that they were trained on.
 We used these pre-trained language models for refining the fluency of the detected episode clusters and did not observe any concerning results, as it is a low-risk consideration for the domains that we studied.
-Appendix B Additional Details for Episode Partitioning
+9 Acknowledgements
+This work was supported by the National Science Foundation Graduate Research Fellowship.
+This research used the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois.
+DeltaAI is a joint effort of the University of Illinois at Urbana-Champaign and its National Center for Supercomputing Applications.
+Appendix A Experimental Settings
+For implementing EpiMine, we use the following hyperparameters across all datasets: $\delta=25\%$, $sim\_thresh=0.75$.
+All other hyperparameters are set to their respective default values.
+We provide all experimental settings in Appendix A.To determine statistical significance, we check for $\geq\mu-\sigma$.
+For our word representations, we use bert-base-uncased.
+For our sentence representations, we use all-mpnet-base-v2.
+We choose to use Claude-2.1444claude.ai/ for fluent candidate episode estimation due to its strong structured JSON/XML input and output formatting abilities.
+However, this proprietary model can be replaced with any open-source model as EpiMine is model-agnostic.
+We use only one NVIDIA GeForce GTX 1080 for all experiments; for non-API models, we utilize two NVIDIA-RTX A6000s.
+Appendix B Baselines
+We compare against the following methods using the evaluation metrics specified in Appendix C.
+  • K-means Likas et al. (2003): No. of ground-truth episodes is given; clusters segments based on semantic similarity of ST  Reimers and Gurevych (2019) embeddings.
+  • EvMine Zhang et al.
+(2022): Unsupervised framework for key event detection that leverages peak phrases and detects communities using event-indicative features.
+We extend the original document-level method to the segment level for episode detection.
+  • EMiner Jiao et al.
+(2023): Unsupervised event chain mining that performs atomic action clustering.
+For episode detection, we map its final output, a list of events, back to the original sentences from which each event was extracted, treating these sentences as segments.
+To retrieve more episode-associated segments, we use ST Reimers and Gurevych (2019) to select the $k$ most similar segments to each cluster sentence.
+We also include the following full and partial ablations of EpiMine (clusters segments from all articles vs. top $\delta$articles, respectively):
+  • No Confidence: A full ablation, where all input segments are classified based on the episode with max cosine similarity instead of using the confidence score from Equation 2.
+  • No LLM: We take the estimated episode clusters from Section 4.3 that normally would have been inputted into the LLM, and instead use them to compute our episode representations directly.
+These representations are used for our classification step (Section 4.4), run on the full dataset with confidence.
+  • EpiMine-Top: A partial ablation which directly outputs the intermediate episode clusters formed based on the top articles identified in Section 4.3 without inputting them into the LLM-based episode estimation step.
+  • TF-IDF: A partial ablation which replaces the salience and synonym expansion step (Section 4.1) with TF-IDF.
+  • No DC: A partial ablation which replaces the discriminative co-occurrence score (Equation 1) with raw pair frequency.
+Appendix C Evaluation Metrics
+Following a recent work on key event detection Zhang et al.
+(2022), we adapt the $k$-prec, $k$-recall, and $k$-F1 metrics to quantitatively evaluate episode detection performance— specifically, how the model’s top-$k$ segments within each detected episode align with the ground truth episodes.
+Formally, suppose there are $N$ ground truth episodes $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$, each of which is a set of text segments related to its corresponding episode.
+=$\mathcal{E}=\{E_{1},E_{2},\dots,E_{K}\}$ are the model predicted episodes, each of which is a ranked list of segments, and $E_{j,k}$ means the top-$k$ segments within $E_{j}$.
+Then, the $k$-metrics are defined as follows:
+$\begin{split}\text{{k-prec}}&=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{\sum_{E_{j}\in\mathcal{E}}\mathds{1}(|E_{j}|\geq k)}\\ \text{{k-recall}}&=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{N}\\ \text{{k-F1}}&=\frac{2\cdot\text{k-prec}\cdot\text{k-recall}}{\text{k-prec}+\text{k-recall}}\\ \end{split}$
+Appendix D Key Event Corpus Dataset Construction & Annotation
+Given that our task is novel and no large-scale key event-specific news corpus is available for this task where the key events are guaranteed to contain distinguishable episodes, we briefly discuss how we collect the input corpus from online news data.
+Given our set of key events (as listed in Section 5.1), we first scrape the external reference list from their corresponding Wikipedia page and select the news articles that have been published within two months of given key event’s start date (e.g., all articles selected for “January 6 2021 Capitol Attack” would have been published between November 6-March 6).
+This is important as we want to prioritize the news articles which focus on describing the episodes of the key event and their corresponding aspects as opposed to primarily opinions or analyses.
+This allows us to motivate our task as one critical for currently evolving key events which required a more fine-grained episodic timeline.
+Furthermore, it is consequently unlikely for a single article to cover all the episodes and exclusively episodes under a key event.
+Despite this being more challenging, it is acceptable as the goal of our task is to extract only the key event-related episodes, which must be substantiated by multiple documents in either case.
+During the collection process, we targeted selecting a diverse set of key events topics within a theme.
+For instance, we attempted to cover every type of “natural disaster”, including tornados, wildfires, and etc.
+When selecting key events, we leave out those with less than $20$ hyperlinks in the Wikipage and manually inspect at least 20 articles per event in order to ensure quality.
+Table 1 summarizes the statistics for these datasets.
+We also construct a background news corpus of approximately 4,000 long news articles using the New York Times corpus for topic categorization Meng et al.
+(2020).
+For the annotation process, we had each of the four individual annotators (computer science graduate students) manually identify a ground-truth description for each of the episodes under every key event.
+The descriptions and a one-shot annotation demonstration (assigning either an episode ID per segment, or ‘M’/‘X’ if it describes multiple or none) are provided in our automatic annotation prompt (Appendix I).
+The same annotators each manually annotated a subset of segments (25 articles, 300 segments in total) with either their corresponding episode ID or ‘-1’ if the segment was either ‘X’ and ‘M’.
+The human-automatic annotation agreement is shown in Table 6 with three versions of intra-class correlation (ICC) and Cohen’s $\kappa$).
+The Cohen’s $\kappa$ indicates substantial agreement, and the all three versions of ICC indicate good reliability.
+Cohen’s $\kappa$ ICC1k ICC2k ICC3k
+Score 0.614 0.772 0.772 0.773
+Table 6: Agreement scores between the human and LLM annotation of each article segment.
+Appendix E Identifying Salient Terms for Episode Detection
+We define the salience score of a term $w_{i}$ within segment $s$ as the following function, where $freq(w_{i})$ is the number of key event segments that $w_{i}$ is contained in, $N_{bg}$ is the number of news articles in the background corpus we construct (using general New York Times articles), and $bgf(w_{i})$ is the number of background articles that $w_{i}$ is present in.
+$\begin{split}\textbf{Salience}(w_{i})=&\left(1+\log^{2}\left(freq(w_{i})\right)\right)\\ &\times\log\left(\frac{N_{bg}}{bgf(w_{i})}\right)\end{split}$ (3)
+Stop words and infrequent terms ($freq(w_{i})<5$) are assigned a salience score of $-1$.
+A key event’s set of salient terms $T$ is comprised of the terms with a salience score above the mean salience across the entire vocabulary.
+In the case of infrequent synonyms used by a journalist as a stylistic choice (e.g., “demonstrations”, “marches”), we expand $T$ with terms that are similar (cosine-similarity) to their static word representations (average of its contextualized word embeddings across entire key event corpus).
+Appendix F Key Event News Article Pre-Processing
+Given that the expected output for the episode detection task is a cluster of text segments, we first must segment each key event news article.
+We would like to ideally preserve both the primary aspects (e.g., core entities and their actions) and peripheral aspects (e.g., reactions to a core entity’s action) relevant to that episode, which may be helpful for cross-document episode co-reference resolution.
+In order to do this, we utilize the text segmentation method, C99 Choi (2000).
+Furthermore, in order to assist with the cohesiveness of the segment, we employ entity co-reference resolution before performing segmentation, which assists with retaining the context across text segments (“They surrounded the legislative building […]” $\rightarrow$ “The protesters surrounded the legislative building […]”).
+Our core methodology is given these text segments (in their raw form, without co-references resolved) and their source articles as the primary inputs.
+Appendix G Additional Details for Episode Partitioning
 We note that for determining semantic similarity between the terms of two segments, we use both (1) the average cosine similarity between all unordered pairs of terms between segment (i-1) and (i), and (2) the cosine similarity between the average of static term representations in $(i-1)$ and the average of static term representations in $(i)$.
 Furthermore, we filter out any non-salient segments before episode partitioning to avoid any influence of noisy segments (e.g., journalist’s analysis, summary statements, historical comparisons, and other generic noise) on the quality of our episode fragments.
-Finally, following (Wang et al., 2021; Kargupta et al., 2023), we take the harmonic mean of all pairwise discriminative co-occurrence scores instead of a simple average.
+Finally, following Wang et al.
+(2021); Kargupta et al.
+(2023), we take the harmonic mean of all pairwise discriminative co-occurrence scores instead of a simple average.
 This allows us to prioritize the more salient and discriminative terms when determining the episode partitions.
 For instance, if “protesters” consistently occurs throughout the majority of episodes and thus has a low average discriminative co-occurrence, then it is not as informative for episode partitioning.
-Appendix C Claude-2 Prompt & Example for Candidate Episode Estimation
+Appendix H Claude-2 Prompt & Example for Candidate Episode Estimation
 Prompt.
@@
 example_sentences’: [’Hundreds of anti-extradition bill protesters finally broke into the legislature after many hours of attacking the public entrance and ransacked the building, including displaying the colonial Hong Kong flag in the chamber.’, ’Slogans on the wall read: ‘‘Murderous regime’’, and ‘‘There are no rioters only a tyrannical regime.’’]
-Appendix D Evaluation Metrics
-Following a recent work on key event detection (Zhang et al., 2022), we adapt the $k$-prec, $k$-recall, and $k$-F1 to quantitatively evaluate the episode detection results.
-We use these metrics to evaluate how the model output matches the ground truth episodes using the top-$k$ segments within each detected episode.
-Formally, suppose there are $N$ ground truth episodes $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$, each of which is a set of text segments related to its corresponding episode.
-=$\mathcal{E}=\{E_{1},E_{2},\dots,E_{K}\}$ are the model predicted episodes, each of which is a ranked list of segments, and $E_{j,k}$ means the top-$k$ segments within $E_{j}$.
-Then, the $k$-metrics are defined as follows:
-k-prec $=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{\sum_{E_{j}\in\mathcal{E}}\mathds{1}(|E_{j}|\geq k)}$
-k-recall $=\frac{\sum_{G_{i}\in\mathcal{G}}\mathds{1}(\exists E_{j}\in\mathcal{E},E_{j,k}\cap G_{i}\geq\frac{k}{2})}{N}$
-k-F1 $=\frac{2\cdot\text{k-prec}\cdot\text{k-recall}}{\text{k-prec}+\text{k-recall}}$
-Appendix E Key Event Corpus Retrieval
-Given that our task is novel and no large-scale key event-specific news corpus is available for this task where the key events are guaranteed to contain distinguishable episodes, we briefly discuss how we collect the input corpus from online news data, even if it is out of scope for our fundamental methodology.
-Given our set of key events (as listed in Section 5.1), we first scrape the external reference list from their corresponding Wikipedia page and select the news articles that have been published within two months of given key event’s start date (e.g., all articles selected for “January 6 2021 Capitol Attack” would have been published between November 6-March 6).
-This is important as we want to prioritize the news articles which focus on describing the episodes of the key event and their corresponding aspects as opposed to primarily opinions or analyses.
-This allows us to motivate our task as one critical for currently evolving key events which required a more fine-grained episodic timeline.
-Furthermore, it is consequently unlikely for a single article to cover all of the episodes and exclusively episodes under a key event.
-Despite this being more challenging, it is acceptable as the goal of our task is to extract only the key event-related episodes, which must be substantiated by multiple documents in either case.
-We list all of our selective themes and their corresponding key events included in the dataset that we construct:
-  • Terrorism and attacks: 2021 Atlanta spa shootings; 2014 Montgomery County Shootings; 2021 Indianapolis FedEx shooting; 2022 Cincinnati FBI field office attack; 2019 Jersey City shooting; 2019 Naval Air Station Pensacola shooting; 2022 Greenwood Park Mall shooting, 2018 Capital Gazette shooting; 2021 Collierville Kroger shooting; 2019 Kyoto Animation arson attack
-  • Natural disasters: 2023 Tornado outbreak sequence; 2023 Hawaii Wildfires; 2021 Western Kentucky tornado; 2017 Mocoa landslide; 2010 Haiti earthquake; 2021 Henan floods; 2019 Nyonoksa radiation accident; 2022 North American winter storm; 2011 Fukushima nuclear accident
-  • Political Events: 2020 Kyrgyz Revolution, 2019 Storming of the Hong Kong Legislative Council Complex; 2019 Siege of the Hong Kong Polytechnic University; 2017 Zimbabwean coup; 2018 Italian government formation; 2021 January 6 United States Capitol attack; 2018 Thai Cave Rescue Operation; 2018 Armenian Revolution; 2017 Lebanon–Saudi Arabia dispute; 2013 Tunisian political crisis
-Appendix F Claude-2 Prompt for Dataset Annotation
+Appendix I Claude-2 Prompt for Dataset Annotation
 We automatically annotate our dataset using Claude-2.1 using the prompt below (before an additional human-verification stage):